### v2

# From Specific-MLLMs to Omni-MLLMs: A Survey on MLLMs Aligned with Multi-modalities

Abstract
To tackle complex tasks in real-world scenarios, more researchers are focusing on Omni-MLLMs, which aim to achieve omni-modal understanding and generation. Beyond the constraints of any specific non-linguistic modality, Omni-MLLMs map various non-linguistic modalities into the embedding space of LLMs and enable the interaction and understanding of arbitrary combinations of modalities within a single model. In this paper, we systematically investigate relevant research and provide a comprehensive survey of Omni-MLLMs. Specifically, we first explain the four core components of Omni-MLLMs for unified multi-modal modeling with a meticulous taxonomy that offers novel perspectives. Then, we introduce the effective integration achieved through two-stage training and discuss the corresponding datasets as well as evaluation. Furthermore, we summarize the main challenges of current Omni-MLLMs and outline future directions. We hope this paper serves as an introduction for beginners and promotes the advancement of related research. Resources will be made public.
From Specific-MLLMs to Omni-MLLMs:
A Survey on MLLMs Aligned with Multi-modalities
Shixin Jiang1, Jiafeng Liang1, Jiyuan Wang1, Xuan Dong1, Heng Chang2 Weijiang Yu2, Jinhua Du2, Ming Liu1,3thanks: Corresponding author, Bing Qin1,3 1Harbin Institute of Technology, Harbin, China 2Huawei Inc., Shenzhen, China 3Peng Cheng Laboratory, Shenzhen, China {sxjiang, jfliang, jywang, mliu, qinb}@ir.hit.edu.cn
1 Introduction
The remarkable performance of continuously evolving Multi-modal Large Language Models (MLLMs) has pointed to a possible direction for achieving general artificial intelligence (Bubeck et al., 2023; OpenAI, 2023b). MLLMs extend Large Language Models (LLMs) by integrating them with pre-trained models tailored to specific modalities, such as Vision-MLLMs (Liu et al., 2023c; Wang et al., 2024b; Sun et al., 2024c), Audio-MLLMs (Zhang et al., 2023a; Chu et al., 2023), and 3D-MLLMs (Xu et al., 2024b). However, these modality-specific MLLMs (Specific-MLLMs) are insufficient to tackle complex tasks in real-world scenarios that simultaneously involve multiple modalities. Therefore, efforts are being made to expand the range of modalities for understanding and generation, giving rise to the omni-modality MLLMs (Omni-MLLMs).
Figure 1: The timeline of representative Omni-MLLMs.
By integrating multiple pre-trained models of more non-linguistic modalities (Radford et al., 2021, 2023; Xue et al., 2024b; Rombach et al., 2022; Liu et al., 2023b), Omni-MLLMs expand the modalities for understanding and even generation based on Specific-MLLMs. Omni-MLLMs leverage the emergent capabilities of LLMs to treat various non-linguistic modalities as different foreign languages, enabling the interaction and understanding of information across different modalities within a unified space (Chen et al., 2023a; Panagopoulou et al., 2024). Compared to Specific-MLLMs, Omni-MLLMs can perform multiple uni-modal111 “Uni” and “Cross” refer to the number of non-linguistic modalities involved in the interaction, in contrast to “multimodal reasoning,” traditionally reserved for vision-language tasks Panagopoulou et al. (2024). understanding and generation tasks, as well as cross-modal tasks across two or more non-linguistic modalities, allowing a single model to handle arbitrary combinations of modalities.
A review of the development of Omni-MLLMs reveals that it has been continuously expanding in three directions. On the one hand, the types of modalities processed by Omni-MLLM have been continuously increasing, from X-LLMs that handle vision and audio to X-InstructBLIP (Panagopoulou et al., 2024) which adds 3D modality capabilities, PandaGPT (Su et al., 2023) that incorporates IMU modality, and finally One-LLM (Han et al., 2024a), which processes eight different modalities simultaneously. On the other hand, the ability to interact across modalities of Omni-MLLMs has also expanded, from the joint 3D-Image and Audio-Image cross-modal reasoning capability in ImageBind-LLM (Han et al., 2023) to the cross-modal generation capability of CoDi-2 that leverages interleaved audio and image contexts to generate both audio and images (Tang et al., 2024b). The Omni-MLLM is thus trending towards an “Any-to-Any” model. Besides, the application scenarios of Omni-MLLMs have been broadened, encompassing real-time multimodal speech interaction like Mini-Omni2 and Lyra (Xie and Wu, 2024; Zhong et al., 2024), world simulation like WordGPT (Ge et al., 2024b), multi-sensor autonomous driving like DriveMLM (Wang et al., 2023b), etc. In addition to the open-source models, there are also some closed-source Omni-MLLMs such as GPT-4o (OpenAI, ), Gemini (Reid et al., 2024), and Reka (Ormazabal et al., 2024). The timeline of Omni-MLLMs is shown in Figure 1. Despite the emergence of numerous Omni-MLLMs, there is still a lack of systematic evaluation and analysis.
To fill the gap, we propose this work to conduct a comprehensive and detailed analysis of Omni-MLLMs. We first review the architecture of Omni-MLLMs in four parts (§2). Next, we summarize how Omni-MLLMs expand across multiple modalities through the two-stage training process (§3); then present the training data construction and performance evaluation (§4). Furthermore, we highlight some key challenges and future directions (§5). Finally, we provide a brief summary (§6) and discuss related surveys in the Appendix A.
Our contributions can be summarized as follows: (1) Comprehensive Survey: This is the first comprehensive survey dedicated for Omni-MLLMs; (2) Meticulous taxonomy: We introduce a meticulous taxonomy (shown in Figure 2); (3) Challenges and Future: We outline the challenges of Omni-MLLMs and shed light on future research.
2 Omni-MLLM Architecture
{forest}
forked edges, for tree= grow’=0, draw, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.5em,font=,, where level=2text width=5.5em,font=,, where level=3text width=5.5em,font=,, where level=4text width=5em,font=,, [ Omni-MLLMs, ver [ Continuous
Encoding [ Multi-branch
Projection [ Cross-modal Understanding: eP-ALM (Shukor et al., 2023a),X-LLM (Chen et al., 2023a), ChatBridge
 (Zhao et al., 2023b), Video-LLaMA (Zhang et al., 2023b), LAMM (Yin et al., 2023b), Macaw-LLM
 (Lyu et al., 2023), BuboGPT (Zhao et al., 2023a), AnyMAL (Moon et al., 2024), FAVOR (Sun et al., 2023a),
Octavius (Chen et al., 2024g), X-InstructBLIP (Panagopoulou et al., 2024), AV-LLM (Shu et al., 2023),
DriveMLM (Wang et al., 2023b), Omni-3D (Li et al., 2024b), MultiPLY (Hong et al., 2024), GroundingGPT
 (Li et al., 2024i), DAMC (Chen et al., 2024a), CAT (Ye et al., 2024b), AVicuna (Tang et al., 2024a),
QaP (Liang et al., 2024), Uni-Moe (Li et al., 2024g), Video-LLaMA2 (Cheng et al., 2024b), Emotion-LLaMA
 (Cheng et al., 2024a), video-SALMONN (Sun et al., 2024a), InternOmni (OpenGV, 2024), EAGLE
 (Koska and Horváth, 2024), Meerkat (Chowdhury et al., 2024), Dolphin (Anonymous, 2024), Med-2E3
 (Shi et al., 2024), Llama-AVSR (Cappellazzo et al., 2024), Baichuan-Omni (Li et al., 2024c), OMCAT
 (Goel et al., 2024), LongVALE-LLM (Geng et al., 2024), Baichuan-Omni1.5 (Li et al., 2025),
OpenOmni (Luo et al., 2025)
Cross-modal Understanding & Generation: NExT-GPT (Wu et al., 2024b), Unified-IO2 (Lu et al., 2024a),
ModaVerse (Wang et al., 2024c), X-VILA (Ye et al., 2024a), REAMO (Zhang et al., 2024b), UnifiedMLLM
 (Li et al., 2024h), VITA (Fu et al., 2024b), Mini-Omni2 (Xie and Wu, 2024), CAD-MLLM (Xu et al., 2024a),
MuMu-LLaMA (Liu et al., 2024b), Lyra (Zhong et al., 2024), VITA-1.5 (Fu et al., 2025),
EmpatheticLLM (Kim et al., 2024b), MiniCPM-o (MiniCPM-o Team, 2025) , leaf, text width=46em ] ] [ Uni-branch
Projection [ Cross-modal Understanding: PandaGPT (Su et al., 2023), ImageBind-LLM (Han et al., 2023), One-LLM
 (Han et al., 2024a), CREMA (Yu et al., 2024c), TVL (Fu et al., 2024c), PathWeave (Yu et al., 2024a)
Cross-modal Understanding & Generation: CoDi-2 (Tang et al., 2024b), WorldGPT (Ge et al., 2024b),
EmpathyEar (Fei et al., 2024), EGMI (Kim et al., 2024a) , leaf, text width=46em ] ] ] [ Discrete
Encoding [Embedding [Cross-modal Understanding & Generation: Teal (Yang et al., 2023b), AnyGPT (Zhan et al., 2024), M3GPT
 (Luo et al., 2024), OccLLaMA (Wei et al., 2024), MIO (Wang et al., 2024g), Gesticulator (Pang et al., 2024),
SOLAMI (Jiang et al., 2024b) , leaf, text width=46em ]] ] [ Hybrid
Encoding [ Projection
&Embedding [ Cross-modal Understanding: SynesLM (Lu et al., 2024b)
Cross-modal Understanding & Generation: LEO (Huang et al., 2024), Ground-Action (Szot et al., 2024a),
EMOVA (Chen et al., 2024c), GMA (Szot et al., 2024b), leaf, text width=46em ]] ] ]
Figure 2: Taxonomy for Omni-MLLMs based on their encoding and alignment methods.
As the extension of Specific-MLLMs, Omni-MLLMs inherit the architecture of encoding, alignment, interaction, and generation and broaden the types of non-linguistic modalities involved. This section introduces the implementation methods and functions of the four components in Omni-MLLM: Multi-modalities Encoding (§2.1), Multi-modalities Alignment (§2.2), Multi-modalities Interaction (§2.3), and Multi-modalities Generation (§2.4). More details about the architecture of Omni-MLLMs are shown in Appendix B.
2.1 Multi-modalities Encoding
Based on the encoding feature spaces of multiple modalities, we categorize the Omni-MLLM encoding methods into three types: 1) continuous encoding, 2) discrete encoding, and 3) hybrid encoding.
2.1.1 Continuous Encoding
Continuous encoding refers to encoding the modality into the continuous feature space. Omni-MLLMs that adopt continuous encoding, such as X-LLM (Chen et al., 2023a) and ChatBridge (Zhao et al., 2023b), often integrate multiple pre-trained uni-modality encoders. These modality-specific encoders encode different modalities $\boldsymbol{X}$ into distinct feature spaces $\mathbb{R}_{x}$ as $\mathbf{F}_{x}$, formulated as:
$\mathbf{F}_{x}=\operatorname{SpecificEncoder}\left(\boldsymbol{X}\right),\ \ \mathbf{F}_{x}\in\mathbb{R}_{x}$ (1)
where SpecificEncoder refers to different modality-specific encoders used in Omni-MLLMs, such as InternVit (Chen et al., 2023g) for encoding visual modality, Whisper (Radford et al., 2023) for encoding auditory modality, ULIP-2 (Xue et al., 2024b) for encoding 3D modality, IMU2CLIP (Moon et al., 2022) for encoding IMU modality, etc.
Besides using heterogeneous encoders for continuous encoding, some Omni-MLLMs (Han et al., 2024a, 2023; Su et al., 2023; Fu et al., 2024c) employ pre-aligned encoders for multiple modalities, encoding different modalities $\boldsymbol{X}$ into the same feature space $\mathbb{R}_{uni}$, as shown in Equation 2.
$\mathbf{F}_{x}=\operatorname{PreAlignEncoder}\left(\boldsymbol{X}\right),\ \ \mathbf{F}_{x}\in\mathbb{R}_{uni}$ (2)
where $\operatorname{PreAlignEncoder}$ refer to encoders that uniformly encode multiple modalities, such as LanguageBind (Zhu et al., 2024a) which uses text as a bridge to align different modalities, and ImageBind Girdhar et al. (2023) which uses images as a bridge to align different modalities.
2.1.2 Discrete Encoding
To better facilitate the seamless integration and generation of new non-linguistic modalities, some Omni-MLLMs, such as AnyGPT (Zhan et al., 2024) and Teal (Yang et al., 2023b), adopt a discrete encoding approach. This method encodes different raw modalities $\boldsymbol{X}$ into the same discrete token space $\mathbb{V}_{uni}$ as $\mathbf{T}_{x}$, formulated as follows:
$\mathbf{T}_{x}=\operatorname{SpecificTokenizer}\left(\boldsymbol{X}\right),\ \ \mathbf{T}_{x}\in\mathbb{V}_{uni}$ (3)
where $\operatorname{SpecificTokenizer}$ refers to different modality-specific tokenizers used in Omni-MLLMs, including the SEED tokenizer (Ge et al., 2024a) based on Vector Quantized Tokenization (VQ), the SpeechTokenizer (Zhang et al., 2023c) based on Residual Vector Quantized Tokenization (RVQ), the AudioTokenizer of Teal (Yang et al., 2023b) based on k-means clustering, and so on.
2.1.3 Hybrid Encoding
Although discrete encoding facilitates the unified processing of different non-linguistic modalities and text compared to continuous encoding, discrete modality tokens often struggle to capture the detailed information inherent in raw continuous modalities (Chen et al., 2024c; Xie and Wu, 2024). Therefore, some Omni-MLLMs combine both encoding approaches instead of a fully discretized manner, choosing different encoding methods for different modalities. For instance, EMOVA (Chen et al., 2024c) uses the discrete S2U tokenizer to encode auditory modalities while employing the continuous encoder InternVit for visual modalities to retain more vision semantic information. Similarly, GroundAction Szot et al. (2024a) encodes visual modalities using the CLIP Vit and action modalities with its trained action tokenizer.
2.2 Multi-modalities Alignment
Omni-MLLMs align the encoded features of various non-linguistic modalities with the embedding space of LLMs. The multi-modality alignment can be categorized into two approaches: 1) projection alignment and 2) embedding alignment.
2.2.1 Projection Alignment
The continuous encoding Omni-MLLMs insert adapters, referred to as projectors, between the encoders and the LLMs. These projectors map the continuously encoded modality features $\mathbf{F}_{x}$ into the text embedding space as $\mathbf{F}_{p}$. As discussed in Section 2.1.1, $\mathbf{F}_{x}$ may either reside in distinct feature spaces $\mathbb{R}_{x}$ or share the same feature space $\mathbb{R}_{uni}$. For the former, multiple projectors are typically employed to align the $\mathbf{F}_{x}$ of each modality into $\mathbb{R}_{t}$ as $\mathbf{F}_{p}$ independently, addressing dimensional mismatch and feature misalignment across modalities Ye et al. (2024a); Lyu et al. (2023); Moon et al. (2024), formulated as follows:
$\mathbf{F}_{p}=\operatorname{SpecificProjector}(\mathbf{F}_{x}),\ \mathbf{F}_{p}\in\mathbb{R}_{t}$ (4)
where $\operatorname{SpecificProjector}$ refers to the modality-specific projector corresponding to different modalities, called multi-branch projection.
For the latter case, besides the multi-branch approach, Omni-MLLMs like PandaGPT (Su et al., 2023) and WorldGPT (Ge et al., 2024b) adopt a shared projector to achieve unified alignment across modalities to reduce the parameters of multiple projectors, as shown in Equation 5.
$\mathbf{F}_{p}=\operatorname{UnifiedProjector}(\mathbf{F}_{x}),\ \mathbf{F}_{p}\in\mathbb{R}_{t}$ (5)
where $\operatorname{UnifiedProjector}$ refers to the unified projector used to align multiple modalities, a design known as the uni-branch projection. A comparison of the two approaches is illustrated in Figure 3.
In terms of the specific implementation of the projector, the most straightforward approach is to use a multi-layer perceptron (MLP) or a single linear layer (Wu et al., 2024b; Cheng et al., 2024b; OpenGV, 2024). Alternatively, attention mechanisms can be employed to compress the encoded information of non-linguistic modalities. This includes cross-attention-based methods like Q-Former (Panagopoulou et al., 2024; Chen et al., 2023a) and Perceiver (Zhao et al., 2023b; Liang et al., 2024), as well as self-attention-based methods such as UPM in OneLLM (Han et al., 2024a). Additionally, BaiChuan-Omni (Li et al., 2024c) and EMOVA (Chen et al., 2024c) incorporate CNNs to compress the projected features, thereby achieving locality preservation (Cha et al., 2024).
It is also worth noting that in multi-branch Omni-MLLMs, different branches may utilize distinct implementations to better accommodate the unique characteristics of each modality (Li et al., 2024i). For example, Uni-MoE Li et al. (2024g) uses a linear projection for the visual modality and a Q-Former for the auditory modality. Meanwhile, uni-branch Omni-MLLMs, when using an attention-based projector, typically design multiple modality-specific learnable vectors to extract key information from various non-linguistic modalities (Yu et al., 2024a; Han et al., 2024a; Yu et al., 2024c).
Figure 3: The three combinations of encoding and alignment in Omni-MLLM are based on different encoding spaces and alignment structures.
2.2.2 Embedding Alignment
As for discrete encoding Omni-MLLMs, the features of non-linguistic modalities are represented as quantized codes, which reside in the same discrete space $\mathbb{V}_{uni}$ as text tokens. Therefore, new modality-specific discrete tokens $\mathbf{T}_{x}$ are embedded into the continuous feature space $\mathbb{R}_{t}$ by modifying the vocabulary of LLMs and the corresponding embeddings layer, as shown in Equation 6.
$\mathbf{F}_{p}=\operatorname{Embedding}(\mathbf{T}_{x}),\ \mathbf{F}_{p}\in\mathbb{R}_{t}$ (6)
where $\operatorname{Embedding}$ refers to the unified embedding layer corresponding to different modalities, which is typically achieved by adding discrete codebooks from various modalities to the vocabulary and expanding the embedding layer of LLMs (Zhan et al., 2024; Yang et al., 2023b; Wei et al., 2024). For instance, AnyGPT extends the vocabulary of the LLaMA-2 by incorporating 17,408 codes across three modalities—image, speech, and music (Zhan et al., 2024). Besides, some works like Ground-Action (Szot et al., 2024a) and LEO (Huang et al., 2024) overwrite infrequently used tokens in the original vocabulary for alignment, as they extend a smaller set of modality-specific discrete tokens.
Additionally, for hybrid encoding models, alignment is achieved by simultaneously employing both the projection method and the embedding method Chen et al. (2024c); Szot et al. (2024b).
Figure 4: Three generation methods in Omni-MLLMs are implemented based on different output spaces of LLM and the corresponding generative models.
2.3 Multi-modalities Interaction
Omni-MLLMs utilize transformer-based LLMs to facilitate information interaction between different modalities within a unified feature space $\mathbb{R}_{t}$. Commonly used LLMs include the LLaMA series (Touvron et al., 2023), the Qwen series (Bai et al., 2023), and others (Cai et al., 2024; Zeng et al., 2024).
For interaction, most Omni-MLLMs Chen et al. (2023a); Ye et al. (2024a); OpenGV (2024); Li et al. (2025) concatenate aligned non-linguistic modality features $\mathbf{F}_{p}$ with textual features $\mathbf{F}_{t}$ at the input level, enabling interaction in a progressive and layer-by-layer manner. Meanwhile, some works, such as ImageBind-LLM (Han et al., 2023) and TVL-LLaMA (Fu et al., 2024c), insert $\mathbf{F}_{p}$ into specific layers or all layers of the LLMs to mitigate the loss of modality information (Shukor et al., 2023a).
In terms of the number of modalities involved in interactions, compared to Specific-MLLMs that are limited to dual-modal interactions between a single non-linguistic modality and text (Liu et al., 2023c; Xu et al., 2024b), Omni-MLLMs not only support multiple dual-modal interactions but also enable omni-multimodal interactions involving more than two non-linguistic modalities (Zhao et al., 2023b; Wang et al., 2024g). For example, X-InstructBLIP (Panagopoulou et al., 2024) enables dual-modal interactions such as vision-text, audio-text, and 3D-text, as well as omni-modal interactions like vision-audio-text and 3D-vision-text, showcasing the ability of Omni-MLLMs to handle arbitrary combinations of modalities.
2.4 Multi-modalities Generation
Omni-MLLMs can output text while also generating non-linguistic modalities by integrating different generation models. As shown in Figure 4, we categorize multi-modalities generation into three types: text-based generation, representation-based generation, and modality-token-based generation.
Text-based
This approach directly utilizes the discrete text output from the LLM to invoke Text-to-X generation models (Liu et al., 2023b; Luo et al., 2023c; Brooks et al., 2023) based on the content of the text. For example, VITA (Fu et al., 2024b) employs TTS tools (RVC-Boss, ) to convert the output text into corresponding speech, while ModelVerse (Wang et al., 2024c) and UnifiedMLLM (Li et al., 2024h) use the text to specify the generation model and utilize the corresponding descriptions to generate different modalities.
Modality-Token-based
Works like MiniOmni-2 (Xie and Wu, 2024) and AnyGPT (Zhan et al., 2024) extend the corresponding LLM head with codebooks from different modality tokenizers to generate modality-specific discrete tokens. These tokens are then decoded using the corresponding de-tokenizers (Esser et al., 2021; Yu et al., 2024b; Zeghidour et al., 2022; Dhariwal et al., 2020) to produce various modalities.
Representation-based
To alleviate the potential noise introduced by discrete tokens, works like X-VILA (Ye et al., 2024a) and NextGPT (Wu et al., 2024b) incorporate modality-specific signal tokens into the vocabulary. They then use transformers or MLPs to map the signal token representations into the ones that are understandable to the multimodal decoders, typically off-the-shelf latent-conditioned diffusion models (Rombach et al., 2022; Tang et al., 2023; Xue et al., 2024a; Blattmann et al., 2023), enabling effective generation capabilities.
3 Omni-MLLM Training
To achieve alignment across different vector spaces and improve instruction-following ability under arbitrary modality settings, Omni-MLLMs extend the standard two-stage training pipeline of Specific-MLLMs: multi-modalities alignment pre-training and multi-modalities instruction fine-tuning.
3.1 Multi-modalities Alignment Pre-training
Multi-modalities alignment pre-training involves input alignment training between the feature spaces of different modalities and the embedding space of LLMs on the encoding side, as well as output alignment training between the embedding space and the input spaces of various modality decoders on the decoding side. Input alignment and output alignment can be carried out separately (Wu et al., 2024b) or simultaneously (Ye et al., 2024a).
3.1.1 Input Alignment
Input alignment mainly uses X-Text paired datasets of different modalities and minimizes the text generation loss of the corresponding description text to optimize. In this phase, continuous encoding Omni-MLLMs normally update parameters of projectors, while discrete encoding Omni-MLLMs adjust the parameters of the embedding layer.
In terms of training order of different modalities alignment, multi-branch Omni-MLLM performs separate alignment training for each modality-specific projector, directly aligning each non-linguistic modality with text and using text as a bridge to align different non-linguistic modalities (Zhao et al., 2023b; Panagopoulou et al., 2024). The uni-branch Omni-MLLM, on the other hand, uses the unified projector for different modalities, which may lead to interference in the alignment performance between different modalities. Thus, Han et al. (2024a) employ a progressive alignment strategy to align multiple modalities in a specific order. In contrast, discrete encoding Omni-MLLMs, like AnyGPT (Zhan et al., 2024) and M3GPT (Luo et al., 2024), mix the alignment data from different modalities and perform alignment simultaneously.
Besides, in addition to directly leveraging X-Text paired datasets from different modalities for direct alignment, PandaGPT (Su et al., 2023), ImageBind-LLM (Han et al., 2023), and VideoLLaMA (Zhang et al., 2023b) utilize the pre-aligned modality feature space $\mathbb{V}_{uni}$ to achieve indirect alignment between other non-linguistic modalities and text by training solely on Image-Text data.
3.1.2 Output Alignment
The training of output alignment typically utilizes the same X-text paired dataset as input alignment and adheres to the identical training sequence. Meanwhile, the training objectives for output alignment vary depending on the multi-modality generation methods in Section 2.4. Token-based generative Omni-MLLMs optimize the extended LLM head by minimizing the text generation loss associated with modality-specific discrete tokens (Lu et al., 2024a; Wei et al., 2024). Representation-based generative Omni-MLLMs generally optimize their output projectors by minimizing the composite loss comprising three components (Xie and Wu, 2024; Yang et al., 2023b): 1) the text generation loss of signal tokens; 2) the L2 distance between the output representation and the condition vector of the corresponding decoder, i.e. MSE loss; and 3) the conditional latent denoising loss Rombach et al. (2022). For text-based generative Omni-MLLMs, as there is no additional output structure, the output alignment training is generally not required (Wang et al., 2024c; Li et al., 2024h).
3.2 Multi-modalities Instruction Fine-tuning
The instruction fine-tuning phase aims to enhance generalization capability under arbitrary modalities of Omni-MLLMs (Panagopoulou et al., 2024; Wu et al., 2024b; Ye et al., 2024a). Instruction fine-tuning primarily utilizes instruction-following datasets and computes the text generation loss for the corresponding responses to optimize. For models with generation capabilities, the loss mentioned in section 3.1.2 may also be incorporated. During this phase, Omni-MLLMs further perform full-scale tuning of the LLM parameters (Han et al., 2024a; Cheng et al., 2024b) or use PEFT techniques (Han et al., 2024b), such as LoRA (Hu et al., 2022), for partial tuning (Wang et al., 2024e).
Compared to Specific-MLLMs, Omni-MLLMs not only leverage multiple uni-modal instruction data of different modalities for training but also use cross-modal instruction data to enhance their cross-modal ability (Ye et al., 2024a; Zhan et al., 2024; Li et al., 2024c). In addition to directly mixing different instruction data for training (Panagopoulou et al., 2024; Ye et al., 2024a), some works like Uni-Moe (Li et al., 2024g) and Lyra (Zhong et al., 2024) adopt a multi-step fine-tuning approach, introducing different uni-modal and cross-modal instruction data in a specific order for training to gradually enhance their uni-modal and cross-modal ability.
3.3 Other Train Recipes
In addition to the general training paradigms mentioned in Section 3, some other useful training recipes are also used. (1) Prior knowledge from Specific-MLLMs: Since Specific-MLLMs have already achieved effective alignment in single-modal scenarios, some Omni-MLLMs directly leverage their well-trained projectors to reduce the training overhead during the alignment phase. For example, InstructBLIP Panagopoulou et al. (2024) and X-LLM Chen et al. (2023a) use the Q-former trained by BLIP2 to align the visual modality, while NaviveMC and DAMC Chen et al. (2024a) further leverage projectors from multiple models to handle alignment for visual, audio, and 3D modalities separately; (2) Additional human preference training: Szot et al. (2024b) and Ye et al. (2024b) adopt HF training methods like PPO and ADPO to better align with human preferences; (3) Modalities Blending: During progressive alignment pre-training or multi-step instruction fine-tuning, some works Han et al. (2024a); Li et al. (2024c); Chen et al. (2024c) mix previously trained modality data with the current new modality data for training to prevent catastrophic forgetting.
4 Data Construction and Evaluation
This section summarizes the construction of modality alignment data and instruction data used in the Omni-MLLM training process (§4.1), as well as the evaluation across four different capabilities (§4.2).
4.1 Training Data
Alignment Data
Omni-MLLMs leverage caption datasets from various modalities to construct X-Text paired data for alignment pre-training, such as the WebVid (Bain et al., 2021) for visual modality and the AudioCaps (Kim et al., 2019) for auditory modality. However, for data-scarce modalities like depth maps and thermal maps, large-scale text-paired data is lacking (Zhu et al., 2024a; Girdhar et al., 2023). To address this, synthetic methods that use DPT models Ranftl et al. (2021); Bhat et al. (2023); Xu et al. (2023) or image translation models (Lee et al., 2023) to convert image-text pairs into other modality text pairs are widely employed Han et al. (2024a); Chen et al. (2024c); Jiang et al. (2024c).Moreover, interleaved datasets (Zhu et al., 2023a) are used for alignment pre-training in some works (Tang et al., 2024b) to enhance the contextual understanding capability.
Instruction Data
Omni-MLLMs not only leverage uni-modal instruction datasets from Specific-MLLMs, but also construct cross-modal instruction data through diverse methods as follows.
(1) Template-based Construction: Most works Sun et al. (2023a); Zhao et al. (2023b); Zhang et al. (2024b) utilize cross-modal downstream datasets (Sanabria et al., 2018; Chen et al., 2020b) combined with predefined templates to construct cross-modal instructions; (2) GPT Generation: Following the paradigm of LLaVA (Liu et al., 2023c), some Omni-MLLMs Lyu et al. (2023); Zhao et al. (2023b) leverage the labels from the annotated dataset (Lin et al., 2014; Bain et al., 2021) or use pre-trained models like SAM (Chen et al., 2023c) and GRIT (Wu et al., 2024a) to extract meta-information (e.g., captions and object categories) of different modalities. Then they employ powerful LLMs (OpenAI, 2023b, a) to generate cross-modal instructions based on the obtained meta-information; (3) T2X Generation: Li et al. (2024g) use TTS tools to convert the Image-Text2Text uni-modal instructions from LLaVA-v1.5 (Liu et al., 2024a) into Image-Speech-Text2Text cross-modal instructions. AnyGPT (Zhan et al., 2024) and NextGPT (Wu et al., 2024b) leverage Text2X models such as DALL-E-3 (Shi et al., 2020) and MusicGen (Copet et al., 2023) to convert the GPT-generated pure text instructions into Xs2Xs cross-modal instructions. Details about training data are shown in Appendix C.1
4.2 Benchmark
We provide a brief overview of the benchmarks used to evaluate Omni-MLLMs. The statistics of the benchmarks are shown in Appendix C.2.
Uni-modal Understanding
Uni-modal understanding assesses the ability of Omni-MLLMs to comprehend and reason on different non-linguistic modalities, including downstream X-Text2Text datasets such as X-Caption Plummer et al. (2015); Xu et al. (2016), X-QA Goyal et al. (2017); Xu et al. (2017a), and X-Classification Deitke et al. (2023), as well as comprehensive multi-task benchmarks Liu et al. (2024d); Fu et al. (2024a, 2023).
Uni-modal Generation
Uni-modal generation aims to evaluate the ability of Omni-MLLMs to generate a single non-linguistic modality, including the Text2X generation task (Kim et al., 2019; Ruiz et al., 2023) and the Text-X2Text editing task Veaux et al. (2017); Perazzi et al. (2016).
Cross-modal Understanding
Cross-modal understanding evaluates the ability of Omni-MLLMs to jointly comprehend and reason across multiple non-linguistic modalities like Image-Speech-Text2Text (Li et al., 2024g; OpenGV, 2024), Video-Audio-Text2Text (Li et al., 2022a, b), and Image-3D-Text2Text (Panagopoulou et al., 2024).
Cross-modal Generation
Cross-modal generation further evaluates the ability of Omni-MLLMs to generate non-linguistic modalities in conjunction with other non-linguistic modality inputs. For example, the Xs-Text2X benchmark proposed by X-VILA (Ye et al., 2024a) includes tasks such as Image-Text2Audio and Image-Audio-Text2Video.
5 Challenges and Future Directions
Despite Omni-MLLMs having showcased remarkable performance on numerous tasks, there are still some challenges that necessitate further research.
5.1 Expansion of modalities
Most Omni-MLLMs can only process 2-3 types of non-linguistic modalities, and they still face several challenges when expanding more modalities.
Training efficiency
The common method that introduces new modalities through additional alignment pre-training and instruction fine-tuning can lead to significant training cost. Leveraging prior knowledge from Specific-MLLMs (Panagopoulou et al., 2024; Chen et al., 2024a) or using pre-aligned encoders for indirect alignment (Han et al., 2023; Su et al., 2023) can help reduce training overhead but may impact cross-modal performance.
Catastrophic forgetting
Expanding new modalities may adjust the shared parameters, potentially causing catastrophic forgetting of previously trained modalities knowledge (Yu et al., 2024a). This issue can be partially mitigated by mixing trained modality data (Han et al., 2024a; Li et al., 2024g) or fine-tuning only the modality-specific parameters Yu et al. (2024a, c), but both approaches make the training process more complex.
Low-resource modalities
Although the data synthesis method in Section 4.1 can help alleviate the lack of text-paired data and instruction data for low-resource modalities (Han et al., 2024a; Jiang et al., 2024c), the absence of real modality may lead to biases in understanding of that modality.
5.2 Cross-modal capabilities
The Omni-MLLMs have achieved promising performance in cross-modal understanding and generation tasks, but there are still some challenges.
Long Context
When the input contains multiple sequence modalities (video, speech…), the length of the multi-modalities token sequence may exceed the context window of LLMs and lead to memory overflow. While methods such as token compressing (Yu et al., 2024c; Li et al., 2024c) or token sampling (Zhan et al., 2024; Zhong et al., 2024) can reduce the number of input tokens, they also result in a decline in cross-modal performance.
Modality Bias
Due to the imbalance in training data volume and the performance disparity among different modality encoders, Omni-MLLMs may tend to pay attention to the dominant modality while neglecting information from other modalities during cross-modal inference. Balancing the data volume across modalities or enhancing the corresponding modality-specific modules could potentially help mitigate this issue (Leng et al., 2024).
Temporal Alignment
When dealing with different modalities that have temporal dependencies, retaining their temporal alignment information is crucial for subsequent cross-modal understanding. Some attempts have been made to preserve the temporal alignment information between audio and video, such as interleaved modality-specific tokens of video and audio (Tang et al., 2024a) and inserting the time-related special tokens into the multi-modalies tokens (Goel et al., 2024).
Data and Benchmark
Although Omni-MLLMs employ various methods in Section 4.1 to generate cross-modal instruction data, there is still significant room for improvement and expansion, including enhancing the diversity of instructions, incorporating longer contextual dialogues, and exploring more diverse modality interaction paradigms. Similarly, cross-modal benchmarks such as OmniBench (Li et al., 2024f) and OmniR (Chen et al., 2024d) still fall short in terms of task richness and instruction diversity when compared to uni-modal benchmarks like MMMU (Yue et al., 2024) and MME (Fu et al., 2023). And the variety of modalities they cover is also relatively limited.
5.3 Application scenarios
The emergence of Omni-MLLM brings new opportunities and possibilities for various applications. (1) Real-time Multi-modalities Interaction: Fu et al. (2025) and Xie and Wu (2024) achieve robust capabilities in both vision and speech understanding, enabling efficient speech-to-speech interactions with vision in real-time. (2) Comprehensive Planning: Wang et al. (2023b) and Szot et al. (2024a) leverage the complementarity across multiple modalities to achieve better path planning and action planning capabilities than planning with vision information only. (3) World Simulator: Ge et al. (2024b) not only understands and generates different modalities but also predicts state transitions for any combination of modalities.
6 Conclusion
In this paper, we provide a comprehensive survey report on Omni-MLLM, offering a comprehensive review of the field. Specifically, we break down Omni-MLLM into four key components and categorize them based on modal encoding and alignment methods. Subsequently, we provide a detailed summary of the training process of Omni-MLLM and the related resources used. We also summarize the current challenges and the future development directions. This paper is the first systematic survey dedicated to Omni-MLLMs. We hope this survey will facilitate further research in this area.
Limitations
This study provides the first comprehensive survey of Omni-MLLMs. Related work, architecture statistics, more details of training and evaluation, as well as other training recipes, can be found in Appendix A,B,C.
We have made our best effort, but there may still be some limitations. On one hand, due to page limitations, we can only provide a concise overview of the core contributions of mainstream Omni-MLLMs, rather than exhaustive technical details. On the other hand, our review primarily covers research from *ACL, NeurIPS, ICLR, ICML, COLING, CVPR, IJCAI, ECCV, and arXiv, and there is a chance that we may have missed some important work published in other venues. We will stay updated with ongoing discussions in the research community and plan to revise our work in the future to include overlooked contributions.
Appendix A Related Survey
With the advent of MLLMs, there are several surveys detailing the current progress of MLLMs. Yin et al. (2023a); Wu et al. (2023); Caffagni et al. (2024) focus on the early Vision-MLLMs, while Çoban et al. (2024) and Ma et al. (2024) respectively summarize the Audio-MLLMs and 3D-MLLMs. Zhang et al. (2024a); Wang et al. (2024a) conduct an investigation into various Specific-MLLMs of different modalities. He et al. (2024); Chen et al. (2024b) discuss the expansion of MLLM’s generative capabilities. Some works discuss MLLMs in specific domains, such as medicine (Xiao et al., 2024), agriculture (Zhu et al., 2024c), and autonomous driving (Cui et al., 2024). Some works highlight some specific tasks such as safety (Fan et al., 2024), hallucination (Bai et al., 2024b), and acceleration (Zhu et al., 2024c). And Li and Lu (2024); Huang and Zhang (2024) focus on the evaluation of MLLM performance.
Distinct from the above-mentioned surveys, this paper focuses on MLLMs that align multiple non-linguistic modalities [^2] Since MLLMs capable of comprehending both video and imagery generally process video as multiple frames and employ a single vision encoder, we categorize them as Specific-MLLMs, i.e. Vision-MLLMs. with LLMs (Omni-MLLMs), enabling cross-modal understanding or cross-modal generation. As the first systematic survey on Omni-MLLMs, we hope our work will serve as an overview of this emerging direction, fostering future research in the field.
Appendix B Details about Omni-MLLMs architures
Table 1 presents the details of the structure of mainstream Omni-MLLMs. We will list some of the pre-trained models used.
B.1 Modality Encoder
Visual Specific-Encoder
Vit (Dosovitskiy et al., 2021), SigCLIP Vit (Zhai et al., 2023), CLIP Vit (Radford et al., 2021), EVA CLIP Vit (Sun et al., 2023e), InternVit (Chen et al., 2023g), DINOv2 Vit (Oquab et al., 2024), DFNCLIP Vit (Fang et al., 2024a), and OpenCLIP ConvNext (Liu et al., 2022b) encode images to obtain continuous features. TimeSformer (Bertasius et al., 2021), VideoMAE (Tong et al., 2022), MAE-DFer (Sun et al., 2023c), Omni-VL (Sun et al., 2023d), Video-Swin (Liu et al., 2022a), and Vivit (Arnab et al., 2021) encode videos to obtain continuous features.
Audio Specific-Encoder
AST (Gong et al., 2021), Beats (Chen et al., 2023d), Whisper (Radford et al., 2023), HuBert (Hsu et al., 2021), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), MERT (Li et al., 2024e), and PANN (Kong et al., 2020) encode the audio modality to obtain continuous features.
3D Specific-Encoder
ULIP2 (Xue et al., 2024b), GD-MAE (Yang et al., 2023a), PointEncoder (Xu et al., 2024b), FrozenCLIP (Huang et al., 2022), and M3D-CLIP (Bai et al., 2024a) encode the 3D modality to obtain continuous features.
Pre-align Uni-Encoder
LanguageBind (Zhu et al., 2024a), ImageBind (Girdhar et al., 2023), Meta-Transformers (Zhang et al., 2023d), TVL (Fu et al., 2024c), and SSVTP (Kerr et al., 2023) encode multiple non-linguistic modalities into a unified feature space and obtain continuous features. TVL, LanguageBind, ImageBind, and SSVTP construct modality-specific encoders for different modalities and achieve multi-modalities alignment through indirect alignment. Meta-Transformers design distinct modality-specific patch embeddings and use a shared encoder to encode multiple modalities.
Other Specific-Encoder
IMU2CLIP (Moon et al., 2022) encodes the IMU modality to obtain continuous features. Individual modality-specific encoders from LanguageBind or ImageBind are often used independently as specific encoders.
B.2 Modality Tokenizer
Visual Tokenizer
VQ-GAN (Esser et al., 2021), DALL-E (Ramesh et al., 2021), BEiT-V2 (Peng et al., 2022), MAGVIT-v2 (Yu et al., 2024b), and SEED (Ge et al., 2023) encode the visual modality into discrete visual tokens, which can be decoded back into the original image using the de-tokenizer.
Audio Tokenizer
Jukebox (Dhariwal et al., 2020), SoundStream (Zeghidour et al., 2022), SpeechTokenizer (Zhang et al., 2023c), Encodec (Défossez et al., 2023), and S2U (Chen et al., 2024c) encode the audio modality into discrete audio tokens, which can be decoded back into the audio using the corresponding de-tokenizer.
Other Tokenizer
Scene Tokenizer (Wei et al., 2024) encodes the 3D modality into discrete 3D tokens. LEO (Huang et al., 2024), Ground-Action (Szot et al., 2024a), OccLLaMA (Wei et al., 2024), and GMA (Szot et al., 2024b) perform discrete encoding of the action modality to obtain corresponding action tokens, which can be decoded back into the original action using the corresponding de-tokenizer. M3GPT (Luo et al., 2024), Gesticulator (Pang et al., 2024), and SOLAMI (Jiang et al., 2024b) perform discrete encoding of the motion modality to obtain corresponding motion tokens, which can be decoded back into the original motion using the corresponding de-tokenizer.
B.3 Modality Generation Model
For image generation, Stable Diffusion (Rombach et al., 2022) and Instruct-Pix2Pix (Brooks et al., 2023) are used. Video generation models include Zeroscope (Cerspense, 2023), VideoFusion (Luo et al., 2023c), VideoCrafter (Chen et al., 2023b), and ModelScope (Wang et al., 2023a). For audio generation, models such as AudioLDM (Liu et al., 2023b), SNAC (Siuzdak et al., 2024), LLaMA-Omni’s audio decoder (Fang et al., 2024b), MusicGen (Copet et al., 2023), and TiCodec (Ren et al., 2024) are utilized. Meanwhile, StyleTTS (Li et al., 2022d) and GPT-SoVITS (RVC-Boss, ) are employed for speech generation.
B.4 LLM Backbone
Commonly used LLMs include the T5 series (Raffel et al., 2020), LLaMA series (Touvron et al., 2023), Qwen series (Bai et al., 2023), Internlm series (Cai et al., 2024), Chatglm series (Zeng et al., 2024), OPT series (Zhang et al., 2022b), Mixtral series (Jiang et al., 2024a), Mistral series (Jiang et al., 2023), Phi series (Gunasekar et al., 2023), and Yi series (Young et al., 2024).
Appendix C Details of Training and evaluation
C.1 Details of Training Data
The statistical results of some commonly used alignment datasets and the instruction data of mainstream Omni-MLLMs are shown in Table 2 and Table 3. There is still a lack of alignment data for data-scarcity modalities and cross-modal instruction data.
Model Capabilities Multi-Modalities Encoding Multi-Modalities Alignment Multi-Modalities Interaction Multi-Modalities Generation
Modalities Method Encoding Model Method Projector Vocabulary Method LLM Modalities Method Generation model
eP-ALM Corss-modal Understanding Visual/Audio Continuous Encoding Vit/TimeSformer/AST multi-branch linear injection OPT
VALOR Corss-modal Understanding Visual/Audio Continuous Encoding Vit/VideoSwin/AST multi-branch MLP injection Bert
X-LLM Corss-modal Understanding Visual/Audio Continuous Encoding Vit/Conformer multi-branch Q-former+Linear concatenate ChatGLM
ChatBridge Corss-modal Understanding Visual/Audio Continuous Encoding EVAL CLIP Vit/Beats multi-branch Preceiver concatenate Vicuna
PandaGPT Corss-modal Understanding Visual/Audio/3D/ IMU/thermal Continuous Encoding ImageBind uni-branch linear concatenate Vicuna
VideoLLama Corss-modal Understanding Visual/Audio Continuous Encoding EVA CLIP Vit/ ImageBind-Audio multi-branch Q-former+Linear concatenate Vicuna
LAMM Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/ FrozenCLIP multi-branch MLP concatenate Vicuna
Macaw-LLM Corss-modal Understanding Visual/Audio Continuous Encoding Vit/Whisper multi-branch Cross-Attention concatenate LLaMA
BuboGPT Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/ImageBind-Audio multi-branch Q-former+Linear concatenate LLaMA
Teal Corss-modal Understanding && Generation Visual/Audio Discrete Encoding VQ-GAN/ Whisper+K-means embedding Added Vocabulary concatenate LLaMA Image Modality-Token-based VQGAN detokenizer
ImageBind-LLM Corss-modal Understanding Visual/Audio/3D Continuous Encoding ImageBind uni-branch MLP injection LLaMA
Next-GPT Corss-modal Understanding && Generation Visual/Audio Continuous Encoding ImageBind multi-branch Linear concatenate Vicuna Image/Video/Audio Representation-based StableDiffusion1.5/ Zeroscope/AudioLDM
Any-MAL Corss-modal Understanding Visual/Audio/IMU Continuous Encoding CLIP Vit/CLAP/ IMU2CLIP multi-branch Preceiver injection LLaMA-2
FAVOR Corss-modal Understanding Visual/Audio Continuous Encoding EVA CLIP Vit/Whisper multi-branch Q-former+Linear concatenate Vicuna
Octavius Corss-modal Understanding Visual/3D Continuous Encoding CLIP Vit/Object-As-Scene multi-branch MLP concatenate Vicuna
LEO Corss-modal Understanding && Generation Visual/3D/Action Hybrid Encoding OpenCLIP Convnext/PointNet++/ LEO’s Action tokenizer multi-branch MLP/ Spatial Transformer Overwrite Vocabulary concatenate Vicuna Action Modality-Token-based LEO’s Action detokenizer
CoDi-2 Corss-modal Understanding && Generation Visual/Audio Continuous Encoding ImageBind uni-branch MLP concatenate LLaMA-2 Image/Video/Audio Representation-based StableDiffusion2.1/ Zeroscope/AudioLDM2
X-InstructBLIP Corss-modal Understanding Visual/Audio/3D Continuous Encoding EVA CLIP Vit/Beats/ ULIP2 multi-branch Q-former+Linear concatenate Vicuna
One-LLM Corss-modal Understanding Visual/Audio/3D/IMU/fMRI/ Normal Map Continuous Encoding Meta-transformer uni-branch UPM(self-attention) concatenate LLaMA-2
AV-LLM Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/CLAP multi-branch Linear concatenate Vicuna
DriveMLM Corss-modal Understanding Visual/3D Continuous Encoding EVA CLIP Vit/GD-MAE multi-branch Q-former+Linear concatenate LLaMA
Omni-3D Corss-modal Understanding Visual/3D Continuous Encoding CLIP Vit/PointNet++ multi-branch MLP concatenate LLaMA-2
ModaVerse Corss-modal Understanding && Generation Visual/Audio Continuous Encoding ImageBind multi-branch Linear concatenate Vicuna Image/Video/Audio Text-based StabelDiffusion /AudioLDM/VideoFusion
MultiPLY Corss-modal Understanding Visual/Audio/3D /thermal/touch Continuous Encoding CLIP Vit/CLAP /ConceptGraph multi-branch MLP/Linear concatenate Vicuna
CREMA Corss-modal Understanding Visual/Audio/3D /thermal/touch/optical Continuous Encoding EVA CLIP Vit+Linear/ Beats+Linear/ConceptFusion+Linear uni-branch Q-former+Linear concatenate Flan-T5
GroundingGPT Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/ImageBind-Audio multi-branch Q-former+Linear/MLP concatenate Vicuna
DAMC Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/Beats/PointBert(PointLLM) multi-branch Q-former+Linear/MLP concatenate Vicuna
AnyGPT Corss-modal Understanding && Generation Visual/Audio Diserect Encoding SEED tokenizer/ Speech tokenizer/Encodec embedding Extend Vocabulary concatenate LLaMA-2 Image/Speech/Music SEED de-tokenizer/Speech de-tokenizer/Encodec de-tokenizer
TVL-LLaMA Corss-modal Understanding Visual/Touch Continuous Encoding TVL Encoders uni-branch MLP injection LLaMA
SSVTP-LLaMA Corss-modal Understanding Visual/Touch Continuous Encoding SSVTP Encoders uni-branch MLP injection LLaMA
CAT Corss-modal Understanding Visual/Audio Continuous Encoding ImageBind multi-branch Linear concatenate LLaMA-2
AVicuna Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/CLAP multi-branch MLP concatenate Vicuna
WorldGPT Corss-modal Understanding && Generation Visual/Audio Continuous Encoding LanguageBind uni-branch Linear concatenate Vicuna Image/Video/Audio Representation-based Stable Diffusion /AudioLDM/Zeroscope
QaP Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/CLAP multi-branch dot attention+Linear injection DeBERTa-V2-XLarge
Uni-Moe Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/Beats multi-branch Q-former+Linear/MLP concatenate LLaMA
M3GPT Corss-modal Understanding Audio/Motion Diserect Encoding Jukebox tokenizer /M3GPT’s Motion tokenier embedding Extend Vocabulary concatenate T5 Music/Motion Modality-Token-based Jukebox de-tokenizer /M3GPT’s Motion de-tokenizer
X-VILA Corss-modal Understanding Visual/Audio Continuous Encoding ImageBind multi-branch MLP concatenate Vicuna Image/Video/Audio Representation-based Stable Diffusion /AudioLDM/VideoCrafter
REAMO Corss-modal Understanding Visual/Audio Continuous Encoding ImageBind multi-branch Linear concatenate Vicuna
VideoLLaMA2 Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/Beats multi-branch MLP concatenate Mistral
Ground-Action Corss-modal Understanding Visual/Action Hybrid Encoding CLIP Vit /Ground-Action action tokenizer multi-branch Preceiver Overwrite Vocabulary concatenate Vicuna Action Modality-Token-based Ground-Action action de-tokenizer
Emotion-LLaMA Corss-modal Understanding && Generation Visual/Audio Continuous Encoding CLIP Vit /HuBert-Chinese multi-branch MLP concatenate LLaMA-2
EmpathyEar Corss-modal Understanding && Generation Visual/Audio Continuous Encoding ImageBind uni-branch Linear concatenate ChatGLM Video/Audio Text-based StyleTTS2/EAT
video-SALMONN Corss-modal Understanding Visual/Audio Continuous Encoding Vit/Beats multi-branch Q-former+Linear concatenate Vicuna
Meerkat Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/CLAP multi-branch MLP concatenate LLaMA-2
InternOmni Corss-modal Understanding Visual/Audio Continuous Encoding Intern Vit/Whisper multi-branch MLP concatenate InternLM-2.5
SynesLM Corss-modal Understanding Visual/Audio Hybird Encoding SigCLIP Vit /XLSR+K-means multi-branch MLP Extend Vocabulary concatenate OPT
UnifiedMLLM Corss-modal Understanding && Generation Visual/Audio Continuous Encoding CLIP Vit /ImageBind-Audio multi-branch Q-former+Linear concatenate OPT Image/Video/Audio Text-based Instruct-pix2pix/ Auffusion/ModelScope
VITA Corss-modal Understanding && Generation Visual/Audio Continuous Encoding CLIP Vit /VITA’s Audio Encoder multi-branch MLP concatenate Mixtral Speech Text-based GPT-SoVITS
OccLLaMA Corss-modal Understanding && Generation 3D/Action Discrete Encoding OccLLaMA’s 3D tokenizer /OccLLaMA’s Action tokenizer embedding Extend Vocabulary concatenate LLaMA-3.1 Action/3D Modality-Token-based OccLLaMA’s 3D de-tokenizer /OccLLaMA’s Action de-tokenizer
Llama-AVSR Corss-modal Understanding && Generation Visual/Audio Continuous Encoding AV-HuBert /Whisper multi-branch MLP concatenate LLaMA-3.1
MIO Corss-modal Understanding && Generation Visual/Audio Discrete Encoding SEED tokenizer /Speech tokenizer embedding Extend Vocabulary concatenate Yi Image/Speech Modality-Token-based SEED de-tokenizer /Speech de-tokenizer
EMOVA Corss-modal Understanding && Generation Visual/Audio Hybird Encoding Intern Vit /EMOVA’s S2U tokenizer multi-branch C-Abstractor Extend Vocabulary concatenate LLaMA-3.1 Speech Modality-Token-based EMOVA’s S2U de-tokenizer
LLM Gesticulator Corss-modal Understanding && Generation Audio/Motion Diserect Encoding MotionRVQ tokenizer /Encodec tokenizer embedding Extend Vocabulary concatenate Qwen-1.5 Audio/Motion Modality-Token-based MotionRVQ de-tokenizer /Encodec de-tokenizer
Baichuan-Omni Corss-modal Understanding Audio/Motion Continuous Encoding SigCLIP /Whisper multi-branch CNN+MLP/Conv-GMLP concatenate
EGMI Corss-modal Understanding && Generation Audio/Motion Continuous Encoding ImageBind uni-branch Linear concatenate Vicuna Image/Audio Representation-based StableDiffusion /AudioLDM
Dolphin Corss-modal Understanding && Generation Visual/Audio Continuous Encoding CLIP Vit/ImageBind-Audio multi-branch MLP concatenate Vicuna
Mini-Omni2 Corss-modal Understanding && Generation Visual/Audio Continuous Encoding CLIP Vit/Whisper multi-branch MLP concatenate Qwen2 Audio Modality-Token-based SNAC de-tokenizer
OMCAT Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/ImageBind-Audio multi-branch Q-former+transformer concatenate Vicuna
PathWeave Corss-modal Understanding Visual/Whisper Continuous Encoding EVA CLIP Vit/Beats /ULIP2 uni-branch Q-former+Linear concatenate Vicuna
CAD-MLLM Corss-modal Understanding Visual/3D Continuous Encoding DINO v2 /Michelangelo multi-branch Preceiver+Linear/Linear concatenate Vicuna
EAGLE Corss-modal Understanding Visual/Audio Continuous Encoding ImageBind multi-branch MLP concatenate LLaMA-2 Image/Video/Audio Text-based StableDiffusion /AudioLDM/Zeroscope
Spider Corss-modal Understanding && Generation Visual/Audio Continuous Encoding ImageBind multi-branch MLP concatenate LLaMA-2 Image/Video/Audio Text-based StableDiffusion /AudioLDM/Zeroscope
Med-2E3 Corss-modal Understanding Visual/3D Continuous Encoding SigCLIP Vit/M3D-CLIP multi-branch Q-former+Linear/MLP concatenate Phi
LongVALE-LLM Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/Beats/Whisper multi-branch MLP concatenate Vicuna
SOLAMI Corss-modal Understanding Audio/Motion Diserect Encoding SpeechTokenizer /SOLAMI’s MotionTokenizer embedding Extend Vocabulary concatenate Vicuna Speech/Motion Modality-Token-based Speech de-tokenizer /SOLAMI’s Motion de-tokenizer
MuMu-LLaMA Corss-modal Understanding && Generation Visual/Audio Continuous Encoding Vit/ViViT/MERT multi-branch Conv+MLP/Conv+Rnn+MLP injection LLaMA-2 Music Representation-based MusicGen
GMA Corss-modal Understanding && Generation Visual/Action Hybird Encoding SigCLIP Vit /GMA’s Action tokenizer multi-branch MLP Overwrite Vocabulary concatenate Qwen-2 Action Modality-Token-based GMA’s Action de-tokenizer
Lyra Corss-modal Understanding && Generation Visual/Audio Continuous Encoding DFNCLIP Vit/Whipser multi-branch MLP concatenate Qwen-2 Audio Representation-based LLaMA-Omni’s audio decoder
VITA-1.5 Corss-modal Understanding && Generation Visual/Audio Continuous Encoding InternVit/VITA’s Audio Encoder multi-branch MLP/CNN+MLP concatenate Mixtral Speech Representation-based TiCodec decoder
EmpatheticLLM Corss-modal Understanding Visual/Audio Continuous Encoding CLIP Vit/Whisper multi-branch Q-former+Linear concatenate Qwen2.5
Table 1: The architectures of mainstream OmniMLLMs. The architectures of 70 Omni-MLLMs are displayed by encoding, alignment, interaction, and generation.
Name Type Modality #Sample
MSCOCO (Lin et al., 2014) X-Text Image,Text 620K
Visual Genome (Krishna et al., 2017b) X-Text Image,Text 4.5M
Flickr30k (Plummer et al., 2015) X-Text Image,Text 158K
SBU (Ordonez et al., 2011) X-Text Image,Text 1M
DCI (Urbanek et al., 2024) X-Text Image,Text 7.8K
BLIP-Capfilt (Li et al., 2022c) X-Text Image,Text 129M
AI Challenger captions (Wu et al., 2017) X-Text Image,Text 1.5M
Wukong Captions (Gu et al., 2022) X-Text Image,Text 101M
CC12M (Changpinyo et al., 2021) X-Text Image,Text 12.4M
CC3M (Sharma et al., 2018) X-Text Image,Text 3.3M
LAION-5B (Schuhmann et al., 2022) X-Text Image,Text 5.9B
Redcaps (Desai et al., 2021) X-Text Image,Text 12M
LAION-COCO (laioncoco, 2022b(bSchuhmann et al.)Schuhmann, Köpf, Vencu, Coombes, and Beaumont) X-Text Image,Text 600M
LAION-CAT (Radenovic et al., 2023) X-Text Image,Text 440M
LAION-AESTHETICS (laion-aesthetics, 2022b(aSchuhmann et al.)Schuhmann, Köpf, Vencu, Coombes, and Beaumont) X-Text Image,Text 120M
ShareGPT4V (Chen et al., 2024e) X-Text Image,Text 1.2M
LAION-115M (Schuhmann et al., 2021a) X-Text Image,Text 115M
Journeydb (Sun et al., 2023b) X-Text Image,Text 4.4M
Multimodal c4 (Zhu et al., 2023b) X-Text-X Image,Text 43.3M
OBELICS (Laurençon et al., 2023) X-Text-X Image,Text 141M
Panda-70M (Chen et al., 2024f) X-Text Video,Text 70M
Webvid2M (Bain et al., 2021) X-Text Video,Text 2M
Valley-Pretrain-703k (Luo et al., 2023a) X-Text Video,Text 703K
Webvid10M (Bain et al., 2021) X-Text Video,Text 10M
YT-Temporal (Zellers et al., 2022) X-Text Video,Text 180M
ActivityNet Captions (Krishna et al., 2017a) X-Text Video,Text 100K
InterVid (Wang et al., 2024d) X-Text Video,Text 10M
MSRVTT (Xu et al., 2016) X-Text Video,Text 200K
ShareGemini (Share, 2024) X-Text Video,Text 530K
AudioSet Gemmeke et al. (2017) X-Text Audio,Text 2.1M
Clotho (Drossos et al., 2020) X-Text Audio,Text 5k
Auto-ACD (Sun et al., 2024b) X-Text Audio,Text 1.5M
AudioCap (Kim et al., 2019) X-Text Audio,Text 46k
WavCaps (Mei et al., 2024) X-Text Audio,Text 403K
AISHELL-1 (Bu et al., 2017) X-Text Audio,Text 128K
AISHELL-2 (Du et al., 2018) X-Text Audio,Text 1M
Gigaspeech (Chen et al., 2021) X-Text Speech,Text
Common Voice (Ardila et al., 2020) X-Text Speech,Text
MLS (Pratap et al., 2020) X-Text Speech,Text
Music caption (Zhan et al., 2024) X-Text Music,Text 100M
Cap3D (Luo et al., 2023b) X-Text 3D,Text 1M
Objaverse (Deitke et al., 2023) X-Text 3D,Text 800K
ScanRefer (Chen et al., 2020a) X-Text 3D,Text 51.5K
Normal Caption (Han et al., 2024a) X-Text Normal,Text 0.5M
Depth Caption (Han et al., 2024a) X-Text Depth,Text 0.5M
NSD (Allen et al., 2022) X-Text fMRI,Text 9K
Ego4d (Grauman et al., 2022) X-Text Video,IMU,Text 528k
PU-VALOR (Tang et al., 2024a) X-Y-Text Video,Audio,Text 114K
VALOR (Chen et al., 2023e) X-Y-Text Video,Audio,Text 16k
VAST (Chen et al., 2023f) X-Y-Text Video,Audio,Text 414k
VIDAL (Zhu et al., 2024a) X-Y-Text Video, Thermal, Depth, Audio 10M
TVL (Fu et al., 2024c) X-Y-Text Image,Touch,Text 44K
M3D-Cap (Bai et al., 2024a) X-Y-Text Image,3D,Text 115K
Table 2: The statistics for alignment datasets in Omni-MLLMs, including single non-linguistic modality text pairing data (X-Text), multiple non-linguistic modalities text pairing data (X-Text-Y), and single non-linguistic modality text interleaved data (X-Text-X).
Name Source Task Modality Construction Method #Sample
XLLM’s SFT Chen et al. (2023a) MiniGPT-4, AISHELL-2, VSDial-CN, ActivityNet Caps Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 10k
ChatBridge’s SFT (Zhao et al., 2023b) MSRVTT, AudioCaps, VQAv2, VG-QA… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, GPT generation 4.4M+209k
Macaw-LLM (Lyu et al., 2023) MSCOCO, Charades, AVSD, VG-QA… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text GPT generation 69K+50K
BuboGPT’s SFT LLaVA, Clotho, VGGSS Uni-Modal Understanding,Cross-Modal Understanding Image,Audio,Text Template Instructionalization, GPT generation 196K
NextGPT’s SFT (Wu et al., 2024b) WebVid, CC3M, AudioCap, Youtube… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT generation+retrieval, T2X generation 20K
AnyMal’s SFT (Moon et al., 2024) Uni-Modal Understanding Image,Video,Audio,Text Manual Annotation, GPT generation 210K
FAVOR’s SFT (Sun et al., 2023a) LLaVA, MSCOCO, Ego4D, LibriSpeech… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, GPT generation
LEO’s SFT (Huang et al., 2024) ScanQA, SQA3D, 3RScan, CLIPort… Uni-Modal Understanding,Cross-Modal Understanding Image,3D,Text,Action Template Instructionalization, GPT generation 220k
CoDi-2’s SFT (Tang et al., 2024b) MIMIC-IT, LAION-400M, AudioSet, Webvid… Uni-Modal Understanding,Uni-Modal Generation Image,Audio,Text Template Instructionalization
X-InstructBLIP’s SFT (Panagopoulou et al., 2024) MSCOCO, Clotho, MSVD, Cap3D… Uni-Modal Understanding Image,Video,Audio,3D,Text Template Instructionalization, GPT generation 1.6M
OneLLM’s SFT (Han et al., 2024a) LLaVA-150K, Clotho, Ego4D, NSD… Uni-Modal Understanding Image,Video,Audio,3D,ImU, Depth,fMRI,Normal,Text Template Instructionalization, T2X generation 2M
AVLLM’s SFT (Shu et al., 2023) ACAV100M, VGGSound, WebVid2M, WavCaps… Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text GPT generation 1.4M
Uni-IO2’s SFT (Lu et al., 2024a) CC3M, AudioSet, Webvid3m, Omni3D… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT generation 775m
ModaVerse’s SFT (Wang et al., 2024c) Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation Image,Video,Audio,Text GPT generation 2M
REAMO’s SFT (Zhang et al., 2024b) Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization 10K
GroundingGPT’s SFT (Li et al., 2024i) Flickr30K, VCR, Activitynet Captions, Clotho… Uni-Modal Understanding Image,Video,Audio,Text GPT Instructionalization 1M
AnyGPT’s SFT (Du et al., 2018) Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Audio,Text GPT Instructionalization, T2X generation 208K
CAT’s SFT (Ye et al., 2024b) VGGSound, AVQA, VideoInstruct100K… Cross-Modal Understanding Video,Audio,Text GPT Instructionalization 100K
AVicuna’s SFT (Tang et al., 2024a) UnAV-100, VideoInstruct100K, ActivityNet Captions, DiDeMo Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text Template Instructionalization 49K
M3DBench’SFT (Li et al., 2024b) Scannet, ScanRefer, ShapeNet… Uni-Modal Understanding,Cross-Modal Understanding Image,3D,Text Template Instructionalization,//GPT Instructionalization 320k
Uni-Moe’s SFT (Li et al., 2024g) LLaVA-Instruct-150K, LibriSpeech, VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 874K
X-VILA’s SFT (Ye et al., 2024a) WebVid, ActivityNetCaption, LLaVA-Instruct-150K… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization
EMOVA’s SFT (Chen et al., 2024c) ShareGPT-4o, MSCOCO, LLaVA-Instruct-150K… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Audio,Text Template Instructionalization, GPT Instructionalization, T2X generation 4.4M
VideoLLaMA2’s SFT (Cheng et al., 2024b) AVQA, AVSD, MusicCaps… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Text Template Instructionalization 1.5M
PathWeave’s SFT (Yu et al., 2024a) VQAV2, MSRVTT, Cap3D… Uni-Modal Understanding Image,Video,Audio,3D,Depth Template Instructionalization, T2X generation 23.2M
Spider’s SFT (Lai et al., 2024) AudioCap, CC3M, Webvid… Cross-Modal Understanding,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation
GMA’s SFT (Szot et al., 2024b) Meta-World,CALVIN, Maniskill… Uni-Modal Understanding,Cross-Modal Understanding, Cross-Modal Generation Image,Text,Action Template Instructionalization 2.2M
OCTAVIUS’s SFT (Chen et al., 2024g) MSCOCO,Bamboo, ScanNet… Uni-Modal Understanding Image,3D,Text Template Instructionalization, GPT Generation
Lyra’s SFT (Zhong et al., 2024) Mini-Gemini, Collected Youtube’s Audio Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation Image,Audio,Text GPT Generation, T2X Generation 1.5M
video-SALMONN’s SFT (Sun et al., 2024a) LibriSpeech,AudioCaps, LLaVA-Instruct-150K… Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text Template Instructionalization, T2X Generation
Meerkat’s SFT (Chowdhury et al., 2024) VGG-SS,AVSBench, AVQA,MUSIC-AVQA… Cross-Modal Understanding Video,Audio,Text Template Instructionalization, GPT Generation 3M
VITA’s SFT (Fu et al., 2024b) ShareGPT4V,LLaVA-Instruct-150K, ShareGTP4o,ShareGemini… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text T2X Generation
Baichuan-omni’s SFT (Li et al., 2024c) vFLAN,VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text T2X Generation
LongVALE-LLM’s SFT (Geng et al., 2024) LongVALE Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text GPT Generation 25.4K
UnifiedMLLM’s SFT (Li et al., 2024h) LISA,SmartEdit… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation 100K
Dolphin’s SFT (Anonymous, 2024) AVQA,Flickr-SoundNet, VGGSound,LLP… Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text Template Instructionalization, GPT Generation
Table 3: The statistics for OmniMLLM’s Instruction Data, including the data sources, interaction forms, involved modalities, and construction methods.
C.2 Details of Benchmark
The statistical data of some commonly used benchmarks are shown in Table 4. Existing benchmarks still require improvements in terms of the number of modalities and the forms of modality interaction.
Name Capability Category Modality Specific-Task Metrics
VQA v2 (Goyal et al., 2017) Unimodal Understanding Image,Text QA Acc
GQA (Hudson and Manning, 2019) Unimodal Understanding Image,Text QA Acc
DocVQA (Mathew et al., 2021) Unimodal Understanding Image,Text QA Acc
IconQA (Lu et al., 2021) Unimodal Understanding Image,Text QA Acc
OCR-VQA (Mishra et al., 2019) Unimodal Understanding Image,Text QA Acc
STVQA (Biten et al., 2019) Unimodal Understanding Image,Text QA Acc
VSR (Liu et al., 2023a) Unimodal Understanding Image,Text QA Acc
Hateful Meme (Kiela et al., 2020) Unimodal Understanding Image,Text QA AUC
OKVQA (Marino et al., 2019) Unimodal Understanding Image,Text QA Acc
VizWiz (Gurari et al., 2018) Unimodal Understanding Image,Text QA Acc
TextVQA (Singh et al., 2019) Unimodal Understanding Image,Text QA Acc
nocap (Agrawal et al., 2019) Unimodal Understanding Image,Text Caption CIDER
ScienceQA (Lu et al., 2022) Unimodal Understanding Image,Text QA Acc
MSCOCO Caption (Lin et al., 2014) Unimodal Understanding Image,Text Caption CIDER,BLEU
Flickr Caption (Plummer et al., 2015) Unimodal Understanding Image,Text Caption CIDER
Visual Dialog (Das et al., 2017) Unimodal Understanding Image,Text Dialogue MRR
RefCOCO (Yu et al., 2016) Unimodal Understanding Image,Text Grounding Acc
RefCOCO+ (Yu et al., 2016) Unimodal Understanding Image,Text Grounding Acc
RefCOCOg (Mao et al., 2016) Unimodal Understanding Image,Text Grounding Acc
A-okvqa (Schwenk et al., 2022) Unimodal Understanding Image,Text QA Acc
POPE (Li et al., 2023b) Unimodal Understanding Image,Text Hallucination Acc
IIIT5K (Mishra et al., 2012) Unimodal Understanding Image,Text OCR WAC(word ACC)
IC13 (Karatzas et al., 2013) Unimodal Understanding Image,Text OCR WAC(word ACC)
IC15 (Karatzas et al., 2015) Unimodal Understanding Image,Text OCR WAC(word ACC)
Total-Text (Chng and Chan, 2017) Unimodal Understanding Image,Text OCR WAC(word ACC)
CUTE80 (Risnumawan et al., 2014) Unimodal Understanding Image,Text OCR WAC(word ACC)
SVT (Wang et al., 2011) Unimodal Understanding Image,Text OCR WAC(word ACC)
SVTP (Phan et al., 2013) Unimodal Understanding Image,Text OCR WAC(word ACC)
COCO-Text (Veit et al., 2016) Unimodal Understanding Image,Text OCR WAC(word ACC)
MMB (Liu et al., 2024d) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
MME (Fu et al., 2023) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
LLaVA-Bench (Liu et al., 2023c) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
Mmmu (Yue et al., 2024) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
SEED (Ge et al., 2023) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
MM-Vet (Yu et al., 2024d) Unimodal Understanding Image,Text Comprehensive Benchmark GPT ACC
ActivityNet-QA (Yu et al., 2019) Unimodal Understanding Video,Text QA Acc
MSRVTT-QA (Xu et al., 2016) Unimodal Understanding Video,Text QA Acc
MSVD-QA (Xu et al., 2017a) Unimodal Understanding Video,Text QA Acc
How2QA (Li et al., 2020) Unimodal Understanding Video,Text QA Acc
NExTQA (Xiao et al., 2021) Unimodal Understanding Video,Text QA ACC
STAR (Wu et al., 2021) Unimodal Understanding Video,Text QA Acc
MSVD-Caption (Xu et al., 2017a) Unimodal Understanding Video,Text QA CIDER
VATEX  (Wang et al., 2019) Unimodal Understanding Video,Text Caption CIDER
MSRVTT-Caption (Xu et al., 2016) Unimodal Understanding Video,Text Caption CIDER,BLEU
Video-ChatGPT Benchmark (Maaz et al., 2024) Unimodal Understanding Video,Text Comprehensive Benchmark GPT ACC,GPT Score
Kinetics-400 Unimodal Understanding Video,Text Classification Acc
Perception test (Patraucean et al., 2023) Unimodal Understanding Video,Text Comprehensive Benchmark GPT ACC
EgoSchema (Mangalam et al., 2023) Unimodal Understanding Video,Text Comprehensive Benchmark GPT ACC
Mvbench (Li et al., 2024a) Unimodal Understanding Video,Text Comprehensive Benchmark GPT ACC
VideoMME (Fu et al., 2024a) Unimodal Understanding Video,Text Comprehensive Benchmark GPT ACC
Charades-STA (Sigurdsson et al., 2016) Unimodal Understanding Video,Text Grounding IoU
AudioCaps (Kim et al., 2019) Unimodal Understanding Audio,Text Caption CIDER,SPICE,METEOR,BLEU,SPIDER
ClothoAQA (Lipping et al., 2022) Unimodal Understanding Audio,Text QA Acc
Vocalsound (Gong et al., 2022) Unimodal Understanding Audio,Text QA Acc
Clotho v1 (Drossos et al., 2020) Unimodal Understanding Audio,Text Caption CIDER
Clotho v2 (Drossos et al., 2020) Unimodal Understanding Audio,Text Caption CIDER
ESC50 (Piczak, 2015) Unimodal Understanding Audio,Text Classification Acc
LibriSpeech (Panayotov et al., 2015) Unimodal Understanding Audio,Text ASR WER
AISHELL-2 (Du et al., 2018) Unimodal Understanding Audio,Text ASR WER
Wenetspeech (Zhang et al., 2022a) Unimodal Understanding Audio,Text ASR WER
MusicCap (Agostinelli et al., 2023) Unimodal Understanding Audio,Text Caption CLAP Score
TUT2017 (Mesaros et al., 2016) Unimodal Understanding Audio,Text Classification Acc
EHSL (Li et al., 2024g) Unimodal Understanding Audio,Text QA Acc
Cap3D Caption (Luo et al., 2023b) Unimodal Understanding 3D,Text Caption CIDER
Objaverse Caption (Deitke et al., 2023) Unimodal Understanding 3D,Text Caption METEOR,ROUGE,BLEU
Cap3D QA (Luo et al., 2023b) Unimodal Understanding 3D,Text QA Acc
Objaverse Classification (Deitke et al., 2023) Unimodal Understanding 3D,Text Classification GPT ACC
Modelnet40 (Wu et al., 2015) Unimodal Understanding 3D,Text Classification Acc
ScanRefer (Chen et al., 2020a) Unimodal Understanding 3D,Text Grounding mAP
Nr3D (Achlioptas et al., 2020) Unimodal Understanding 3D,Text Caption BLEU,CIDER,METEOR,ROUGE-L
SQA3D (Ma et al., 2023) Unimodal Understanding 3D,Text QA Acc
ScanQA (Azuma et al., 2022) Unimodal Understanding 3D,Text QA Acc
SUN RGB-D (Song et al., 2015) Unimodal Understanding Depth,Text Classification Acc
NYUv2 (Silberman et al., 2012) Unimodal Understanding Depth,Text Classification Acc
SUN RGB-D_generated Nomral (Han et al., 2024a) Unimodal Understanding Normal,Text Classification Acc
NYUv2_generated Nomral (Han et al., 2024a) Unimodal Understanding Normal,Text Classification Acc
ThermalQA (Yu et al., 2024c) Unimodal Understanding Thermal, Text QA Acc
TochQA (Yu et al., 2024c) Unimodal Understanding Touch, Text QA Acc
Ego4D (Grauman et al., 2022) Unimodal Understanding IMU,Text Caption CIDER,ROUGE
NSD (Allen et al., 2022) Unimodal Understanding fMRI,Depth Map,Text Caption CIDER,ROUGE
MSCOCO (Lin et al., 2014) Unimodal Generation Image,Text TX2X Edit FID,CLIPSIM
MSRVTT (Xu et al., 2016) Unimodal Generation Video,Text T2X Generate CLIPSIM
AudioCaps (Kim et al., 2019) Unimodal Generation Audio,Text T2X Generate,TX2X Edit FAD
DAVIS (Perazzi et al., 2016) Unimodal Generation Video,Text TX2X Edit CLIPSIM
UCF-101 (Soomro et al., 2012) Unimodal Generation Video,Text T2X Generate FID,FVD,IS,CLIPSIM
Evalcrafter (Liu et al., 2024c) Unimodal Generation Video,Text T2X Generate FVD,CLIPSIM
VCTK (Veaux et al., 2017) Unimodal Generation Audio,Text T2X Generate,TX2X Edit WER,MCD
MusicCap (Agostinelli et al., 2023) Unimodal Generation Audio,Text T2X Generate FAD
Dreambench (Ruiz et al., 2023) Unimodal Generation Image,Text T2X Generate CLIP-I,CLIP-T,DINO
MUSIC-AVQA (Li et al., 2022b) Crossmodal Understanding Video,Audio,Text QA Acc
AVSD (AlAmri et al., 2018) Crossmodal Understanding Video,Audio,Text Dialogue CIDER,BLEU
RACE-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
VALOR Caption (Chen et al., 2023e) Crossmodal Understanding Video,Audio,Text Caption CIDER,BLEU
MMBench-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
AVQA (Li et al., 2022a) Crossmodal Understanding Video,Audio,Text QA Acc
MCUB (Chen et al., 2024a) Crossmodal Understanding Image,Video,Audio,3D,Text Comprehensive Benchmark Acc
DisCRn (Panagopoulou et al., 2024) Crossmodal Understanding Image,Video,Audio,3D Comprehensive Benchmark Acc
OmniXR (Chen et al., 2024d) Crossmodal Understanding Image,Video,Audio,Text Comprehensive Benchmark Acc
Curse (Leng et al., 2024) Crossmodal Understanding Image,Video,Audio,Text Hallucination Acc
ISQA (Sun et al., 2023a) Crossmodal Understanding Image,Audio,Text QA Acc
VGGSound (Chen et al., 2020b) Crossmodal Understanding Video,Audio,Text QA Acc
VATEX (Wang et al., 2019) Crossmodal Understanding Video,Audio,Text Caption CIDER
UnAV-100 (Geng et al., 2023) Crossmodal Understanding Video,Audio,Text Ground IoU
LLP (Tian et al., 2020) Crossmodal Understanding Video,Audio,Text Ground IoU
Presentation-QA (Sun et al., 2024a) Crossmodal Understanding Video,Audio,Text QA ACC
LongVALE Caption Crossmodal Understanding Video,Audio,Text Caption CIDER
TVL Benchmark (Fu et al., 2024c) Cross-modal Understanding Touch,Image,Text QA ACC
AVEB (Sun et al., 2023a) Crossmodal Understanding,Unimodal Understanding Image,Video,Audio,Text Comprehensive Benchmark ACC,METEOR,SPIDER,WER
XtoX Benchmark (Ye et al., 2024a) Crossmodal Understanding,Crossmodal Generation Image,Video,Audio,Text Comprehensive Benchmark X-to-X Alignment Score
Table 4: An overview of benchmarks and tasks of Omni-MLLMs, including the abilities being evaluated, the involved modalities, specific tasks, and evaluation metrics.
C.3 Performance of Omni-MLLMs
We statistic the performance of various mainstream Omni-MLLMs in uni-modal understanding, cross-modal understanding, and cross-modal, as shown in Table 5. We also show the performance of several Specific-MLLMs (Lin et al., 2024b; Li et al., 2024a; Chu et al., 2023; Xu et al., 2024b; Sun et al., 2024c; Jin et al., 2024a) on selected tasks for comparison. The results are mainly from corresponding papers (some results are used as baselines in other papers). It is worth noting that due to differences in the size and performance of the pre-trained models, Omni-MLLMs with the same backbone LLM may still not be fairly comparable. Therefore, this table only provides a rough trend of performance.
It can be seen from the table that most Omni-MLLMs still exhibit a significant performance gap in uni-modal understanding tasks compared to Specific-MLLMs. Meanwhile, in uni-modal generation tasks, models like AnyGPT and CoDi-2 have achieved performance close to or even surpassing Specific-MLLMs. Additionally, Omni-MLLMs are capable of performing cross-modal tasks that Specific-MLLMs cannot handle.
Model LLM Uni-Modal Understanding Uni-Modal Generation Cross-Modal Understanding
[…]
Omni-MLLMs
eP-ALM OPT-2.7B 38.4 38.51 54.47 61.86
ChatBridge 13B Vicuna-13B 45.3 82.5 43
PandaGPT Vicuna-13B 46.7 23.7 32.7 26.1 33.7 79.8
Video-LLaMA Vicuna-7B 51.6 29.6 40.8 36.7 36.6 81
Macaw-LLM LLAMA-7B 42.1 25.5 3.84 33.3 36.1 34.3 31.8 78.7
ImageBind-LLM LLaMA-7B 23.49 10.3 31 39.72 54.26
NExT-GPT Vicuna-7B 64.5 58.4 66.7 84.5 58 81.3 10.07 8.67 31.97
AnyMAL 13B LLaMA2-13B 59.6
AnyMAL 70B LLaMA2-70B 64.2 95.9 77.8
X-InstructBLIP 7B Vicuna-7B 51.7 41.3 30.61 82.1 8.96 67.9 15.4 50 28.1
X-InstructBLIP 13B Vicuna-13B 49.2 74.7 53.7 21.7 20.3 52.1 44.5 44.23
OneLLM 7B LLaMA2-7B 56.5 71.6 78.6 60 57.9 44.5 47.6
AV-LLM Vicuna-7B 67.3 53.7 35.5 47.6 52.6 45.2
UIO-2xxl 6.8B 52.2 41.5 79.4 71.5 48.9 13.39 5.89
ModaVerse Vicuna-7b 56.5 79.2 11.24 8.22 30.14
CREMA 7B Mistral-7B 52.6
GroundingGPT Vicuna-7B 67.8 51.6 78.7 63.8
NaiveMC Vicuna-7B 55 53.63 80.7
DAMC Vicuna-7B 60.5 57.32 81.31
AnyGPT LLaMA2-7B
CAT LLaMA2-7B 62.7 48.6 92
AVicuna Vicuna-7B 70.2 59.7 53.1 49.6
Uni-MoE LLaMA-7B 55.6 66.2 69.82 32.6
X-VILA 7B Vicuna-7B 72.9
VideoLLaMA2-7B Mistral-7B 71.7 71.4 57.2 80.9
Meerkat Llama-2-7B-Chat 87.14
InternOmni InternLM-2-Chat-7B 81.7
UnifiedMLLM Vicuna-7B
VITA Mixtral-8x7B 71.8
EMOVA LLaMA-3.1-8B 82.8
BaiChuan-omni-7B 72.2 76.2
OMCAT Vicuna-7B 49.4 73.8 90.2
PathWeave-7B Vicuna-7B 47.8 37.4 64 33.5
Spider Llama-2-7B 81.7 11.23 8.18 30.97
Specifc-MLLMs
VILA-7B LLaMA-2-7B 79.9 74.7 68.9
VideoChat2 Vicuna-7B 70 54.1
Qwen-Audio Qwen-7B 57.9
PointLLM Vicuna-7B 47.5
Emu-13B 52 11.66
Video-LaVIT Llama2-7B 73.2 80.3 67.3 30.12
Table 5: The performance of Omni-MLLMs on different benchmarks. The selected uni-modal understanding benchmarks include Video-Text2Text (Xu et al., 2017a, 2016), Image-Text2Text (Goyal et al., 2017; Plummer et al., 2015; Liu et al., 2024d), Audio-Text2Text (Kim et al., 2019; Lipping et al., 2022), and 3D-Text2Text (Deitke et al., 2023). The chosen uni-modal generation benchmarks include Text2Image (Lin et al., 2014), Text2Video (Xu et al., 2016), and Text2Audio (Kim et al., 2019). The selected cross-modal understanding benchmarks are Image-Audio-Text2Text (Chen et al., 2020b; AlAmri et al., 2018) and Video-Audio-Text2Text (Li et al., 2022b, a).

### v3

@@
 We hope this paper serves as an introduction for beginners and promotes the advancement of related research.
-Resources will be made public.
+Resources have been made publicly available at https://github.com/threegold116/Awesome-Omni-MLLMs.
 From Specific-MLLMs to Omni-MLLMs:
@@
 Encoding [ Multi-branch
-Projection [ Cross-modal Understanding: eP-ALM (Shukor et al., 2023a),X-LLM (Chen et al., 2023a), ChatBridge
+Projection [ Cross-modal Understanding: eP-ALM (Shukor et al., 2023),X-LLM (Chen et al., 2023a), ChatBridge
  (Zhao et al., 2023b), Video-LLaMA (Zhang et al., 2023b), LAMM (Yin et al., 2023b), Macaw-LLM
@@
 DriveMLM (Wang et al., 2023b), Omni-3D (Li et al., 2024b), MultiPLY (Hong et al., 2024), GroundingGPT
- (Li et al., 2024i), DAMC (Chen et al., 2024a), CAT (Ye et al., 2024b), AVicuna (Tang et al., 2024a),
-QaP (Liang et al., 2024), Uni-Moe (Li et al., 2024g), Video-LLaMA2 (Cheng et al., 2024b), Emotion-LLaMA
+ (Li et al., 2024h), DAMC (Chen et al., 2024a), CAT (Ye et al., 2024b), AVicuna (Tang et al., 2024a),
+QaP (Liang et al., 2024), Uni-Moe (Li et al., 2024f), Video-LLaMA2 (Cheng et al., 2024b), Emotion-LLaMA
  (Cheng et al., 2024a), video-SALMONN (Sun et al., 2024a), InternOmni (OpenGV, 2024), EAGLE
@@
 ModaVerse (Wang et al., 2024c), X-VILA (Ye et al., 2024a), REAMO (Zhang et al., 2024b), UnifiedMLLM
- (Li et al., 2024h), VITA (Fu et al., 2024b), Mini-Omni2 (Xie and Wu, 2024), CAD-MLLM (Xu et al., 2024a),
+ (Li et al., 2024g), VITA (Fu et al., 2024b), Mini-Omni2 (Xie and Wu, 2024), CAD-MLLM (Xu et al., 2024a),
 MuMu-LLaMA (Liu et al., 2024b), Lyra (Zhong et al., 2024), VITA-1.5 (Fu et al., 2025),
@@
 Encoding [Embedding [Cross-modal Understanding & Generation: Teal (Yang et al., 2023b), AnyGPT (Zhan et al., 2024), M3GPT
- (Luo et al., 2024), OccLLaMA (Wei et al., 2024), MIO (Wang et al., 2024g), Gesticulator (Pang et al., 2024),
+ (Luo et al., 2024), OccLLaMA (Wei et al., 2024), MIO (Wang et al., 2024f), Gesticulator (Pang et al., 2024),
 SOLAMI (Jiang et al., 2024b) , leaf, text width=46em ]] ] [ Hybrid
@@
 Additionally, BaiChuan-Omni (Li et al., 2024c) and EMOVA (Chen et al., 2024c) incorporate CNNs to compress the projected features, thereby achieving locality preservation (Cha et al., 2024).
-It is also worth noting that in multi-branch Omni-MLLMs, different branches may utilize distinct implementations to better accommodate the unique characteristics of each modality (Li et al., 2024i).
+It is also worth noting that in multi-branch Omni-MLLMs, different branches may utilize distinct implementations to better accommodate the unique characteristics of each modality (Li et al., 2024h).
 For example, Uni-MoE Li et al.
-(2024g) uses a linear projection for the visual modality and a Q-Former for the auditory modality.
+(2024f) uses a linear projection for the visual modality and a Q-Former for the auditory modality.
 Meanwhile, uni-branch Omni-MLLMs, when using an attention-based projector, typically design multiple modality-specific learnable vectors to extract key information from various non-linguistic modalities (Yu et al., 2024a; Han et al., 2024a; Yu et al., 2024c).
@@
 (2025) concatenate aligned non-linguistic modality features $\mathbf{F}_{p}$ with textual features $\mathbf{F}_{t}$ at the input level, enabling interaction in a progressive and layer-by-layer manner.
-Meanwhile, some works, such as ImageBind-LLM (Han et al., 2023) and TVL-LLaMA (Fu et al., 2024c), insert $\mathbf{F}_{p}$ into specific layers or all layers of the LLMs to mitigate the loss of modality information (Shukor et al., 2023a).
-In terms of the number of modalities involved in interactions, compared to Specific-MLLMs that are limited to dual-modal interactions between a single non-linguistic modality and text (Liu et al., 2023c; Xu et al., 2024b), Omni-MLLMs not only support multiple dual-modal interactions but also enable omni-multimodal interactions involving more than two non-linguistic modalities (Zhao et al., 2023b; Wang et al., 2024g).
+Meanwhile, some works, such as ImageBind-LLM (Han et al., 2023) and TVL-LLaMA (Fu et al., 2024c), insert $\mathbf{F}_{p}$ into specific layers or all layers of the LLMs to mitigate the loss of modality information (Shukor et al., 2023).
+In terms of the number of modalities involved in interactions, compared to Specific-MLLMs that are limited to dual-modal interactions between a single non-linguistic modality and text (Liu et al., 2023c; Xu et al., 2024b), Omni-MLLMs not only support multiple dual-modal interactions but also enable omni-multimodal interactions involving more than two non-linguistic modalities (Zhao et al., 2023b; Wang et al., 2024f).
 For example, X-InstructBLIP (Panagopoulou et al., 2024) enables dual-modal interactions such as vision-text, audio-text, and 3D-text, as well as omni-modal interactions like vision-audio-text and 3D-vision-text, showcasing the ability of Omni-MLLMs to handle arbitrary combinations of modalities.
@@
 This approach directly utilizes the discrete text output from the LLM to invoke Text-to-X generation models (Liu et al., 2023b; Luo et al., 2023c; Brooks et al., 2023) based on the content of the text.
-For example, VITA (Fu et al., 2024b) employs TTS tools (RVC-Boss, ) to convert the output text into corresponding speech, while ModelVerse (Wang et al., 2024c) and UnifiedMLLM (Li et al., 2024h) use the text to specify the generation model and utilize the corresponding descriptions to generate different modalities.
+For example, VITA (Fu et al., 2024b) employs TTS tools (RVC-Boss, ) to convert the output text into corresponding speech, while ModelVerse (Wang et al., 2024c) and UnifiedMLLM (Li et al., 2024g) use the text to specify the generation model and utilize the corresponding descriptions to generate different modalities.
 Modality-Token-based
@@
 (2022).
-For text-based generative Omni-MLLMs, as there is no additional output structure, the output alignment training is generally not required (Wang et al., 2024c; Li et al., 2024h).
+For text-based generative Omni-MLLMs, as there is no additional output structure, the output alignment training is generally not required (Wang et al., 2024c; Li et al., 2024g).
 3.2 Multi-modalities Instruction Fine-tuning
@@
 Compared to Specific-MLLMs, Omni-MLLMs not only leverage multiple uni-modal instruction data of different modalities for training but also use cross-modal instruction data to enhance their cross-modal ability (Ye et al., 2024a; Zhan et al., 2024; Li et al., 2024c).
-In addition to directly mixing different instruction data for training (Panagopoulou et al., 2024; Ye et al., 2024a), some works like Uni-Moe (Li et al., 2024g) and Lyra (Zhong et al., 2024) adopt a multi-step fine-tuning approach, introducing different uni-modal and cross-modal instruction data in a specific order for training to gradually enhance their uni-modal and cross-modal ability.
+In addition to directly mixing different instruction data for training (Panagopoulou et al., 2024; Ye et al., 2024a), some works like Uni-Moe (Li et al., 2024f) and Lyra (Zhong et al., 2024) adopt a multi-step fine-tuning approach, introducing different uni-modal and cross-modal instruction data in a specific order for training to gradually enhance their uni-modal and cross-modal ability.
 3.3 Other Train Recipes
@@
 (2024a); Chen et al.
-(2024c); Jiang et al.
-(2024c).Moreover, interleaved datasets (Zhu et al., 2023a) are used for alignment pre-training in some works (Tang et al., 2024b) to enhance the contextual understanding capability.
+(2024c); Zhu et al.
+(2024a).
+Moreover, interleaved datasets (Zhu et al., 2023a) are used for alignment pre-training in some works (Tang et al., 2024b) to enhance the contextual understanding capability.
 Instruction Data
@@
 Then they employ powerful LLMs (OpenAI, 2023b, a) to generate cross-modal instructions based on the obtained meta-information; (3) T2X Generation: Li et al.
-(2024g) use TTS tools to convert the Image-Text2Text uni-modal instructions from LLaVA-v1.5 (Liu et al., 2024a) into Image-Speech-Text2Text cross-modal instructions.
+(2024f) use TTS tools to convert the Image-Text2Text uni-modal instructions from LLaVA-v1.5 (Liu et al., 2024a) into Image-Speech-Text2Text cross-modal instructions.
 AnyGPT (Zhan et al., 2024) and NextGPT (Wu et al., 2024b) leverage Text2X models such as DALL-E-3 (Shi et al., 2020) and MusicGen (Copet et al., 2023) to convert the GPT-generated pure text instructions into Xs2Xs cross-modal instructions.
@@
 (2017); Xu et al.
-(2017a), and X-Classification Deitke et al.
+(2017), and X-Classification Deitke et al.
 (2023), as well as comprehensive multi-task benchmarks Liu et al.
@@
 Cross-modal Understanding
-Cross-modal understanding evaluates the ability of Omni-MLLMs to jointly comprehend and reason across multiple non-linguistic modalities like Image-Speech-Text2Text (Li et al., 2024g; OpenGV, 2024), Video-Audio-Text2Text (Li et al., 2022a, b), and Image-3D-Text2Text (Panagopoulou et al., 2024).
+Cross-modal understanding evaluates the ability of Omni-MLLMs to jointly comprehend and reason across multiple non-linguistic modalities like Image-Speech-Text2Text (Li et al., 2024f; OpenGV, 2024), Video-Audio-Text2Text (Li et al., 2022a, b), and Image-3D-Text2Text (Panagopoulou et al., 2024).
 Cross-modal Generation
@@
 Expanding new modalities may adjust the shared parameters, potentially causing catastrophic forgetting of previously trained modalities knowledge (Yu et al., 2024a).
-This issue can be partially mitigated by mixing trained modality data (Han et al., 2024a; Li et al., 2024g) or fine-tuning only the modality-specific parameters Yu et al.
+This issue can be partially mitigated by mixing trained modality data (Han et al., 2024a; Li et al., 2024f) or fine-tuning only the modality-specific parameters Yu et al.
 (2024a, c), but both approaches make the training process more complex.
 Low-resource modalities
-Although the data synthesis method in Section 4.1 can help alleviate the lack of text-paired data and instruction data for low-resource modalities (Han et al., 2024a; Jiang et al., 2024c), the absence of real modality may lead to biases in understanding of that modality.
+Although the data synthesis method in Section 4.1 can help alleviate the lack of text-paired data and instruction data for low-resource modalities (Han et al., 2024a; infraredllava), the absence of real modality may lead to biases in understanding of that modality.
 5.2 Cross-modal capabilities
@@
 Although Omni-MLLMs employ various methods in Section 4.1 to generate cross-modal instruction data, there is still significant room for improvement and expansion, including enhancing the diversity of instructions, incorporating longer contextual dialogues, and exploring more diverse modality interaction paradigms.
-Similarly, cross-modal benchmarks such as OmniBench (Li et al., 2024f) and OmniR (Chen et al., 2024d) still fall short in terms of task richness and instruction diversity when compared to uni-modal benchmarks like MMMU (Yue et al., 2024) and MME (Fu et al., 2023).
+Similarly, cross-modal benchmarks such as OmniBench (Li et al., 2024e) and OmniR (Chen et al., 2024d) still fall short in terms of task richness and instruction diversity when compared to uni-modal benchmarks like MMMU (Yue et al., 2024) and MME (Fu et al., 2023).
 And the variety of modalities they cover is also relatively limited.
@@
 (2024b) discuss the expansion of MLLM’s generative capabilities.
-Some works discuss MLLMs in specific domains, such as medicine (Xiao et al., 2024), agriculture (Zhu et al., 2024c), and autonomous driving (Cui et al., 2024).
-Some works highlight some specific tasks such as safety (Fan et al., 2024), hallucination (Bai et al., 2024b), and acceleration (Zhu et al., 2024c).
+Some works discuss MLLMs in specific domains, such as medicine (Xiao et al., 2024), agriculture (Zhu et al., 2024b), and autonomous driving (Cui et al., 2024).
+Some works highlight some specific tasks such as safety (Fan et al., 2024), hallucination (Bai et al., 2024b), and acceleration (Zhu et al., 2024b).
 And Li and Lu (2024); Huang and Zhang (2024) focus on the evaluation of MLLM performance.
@@
 Audio Specific-Encoder
-AST (Gong et al., 2021), Beats (Chen et al., 2023d), Whisper (Radford et al., 2023), HuBert (Hsu et al., 2021), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), MERT (Li et al., 2024e), and PANN (Kong et al., 2020) encode the audio modality to obtain continuous features.
+AST (Gong et al., 2021), Beats (Chen et al., 2023d), Whisper (Radford et al., 2023), HuBert (Hsu et al., 2021), CLAP (Elizalde et al., 2023), Conformer (Gulati et al., 2020), MERT (Li et al., 2024d), and PANN (Kong et al., 2020) encode the audio modality to obtain continuous features.
 3D Specific-Encoder
@@
 ShareGPT4V (Chen et al., 2024e) X-Text Image,Text 1.2M
-LAION-115M (Schuhmann et al., 2021a) X-Text Image,Text 115M
+LAION-115M (Schuhmann et al., 2021) X-Text Image,Text 115M
 Journeydb (Sun et al., 2023b) X-Text Image,Text 4.4M
@@
 REAMO’s SFT (Zhang et al., 2024b) Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization 10K
-GroundingGPT’s SFT (Li et al., 2024i) Flickr30K, VCR, Activitynet Captions, Clotho… Uni-Modal Understanding Image,Video,Audio,Text GPT Instructionalization 1M
+GroundingGPT’s SFT (Li et al., 2024h) Flickr30K, VCR, Activitynet Captions, Clotho… Uni-Modal Understanding Image,Video,Audio,Text GPT Instructionalization 1M
 AnyGPT’s SFT (Du et al., 2018) Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Audio,Text GPT Instructionalization, T2X generation 208K
@@
 M3DBench’SFT (Li et al., 2024b) Scannet, ScanRefer, ShapeNet… Uni-Modal Understanding,Cross-Modal Understanding Image,3D,Text Template Instructionalization,//GPT Instructionalization 320k
-Uni-Moe’s SFT (Li et al., 2024g) LLaVA-Instruct-150K, LibriSpeech, VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 874K
+Uni-Moe’s SFT (Li et al., 2024f) LLaVA-Instruct-150K, LibriSpeech, VideoInstruct100K… Uni-Modal Understanding,Cross-Modal Understanding Image,Video,Audio,Text Template Instructionalization, T2X generation 874K
 X-VILA’s SFT (Ye et al., 2024a) WebVid, ActivityNetCaption, LLaVA-Instruct-150K… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization
@@
 LongVALE-LLM’s SFT (Geng et al., 2024) LongVALE Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text GPT Generation 25.4K
-UnifiedMLLM’s SFT (Li et al., 2024h) LISA,SmartEdit… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation 100K
+UnifiedMLLM’s SFT (Li et al., 2024g) LISA,SmartEdit… Uni-Modal Understanding,Cross-Modal Understanding, Uni-Modal Generation,Cross-Modal Generation Image,Video,Audio,Text Template Instructionalization, GPT Generation 100K
 Dolphin’s SFT (Anonymous, 2024) AVQA,Flickr-SoundNet, VGGSound,LLP… Uni-Modal Understanding,Cross-Modal Understanding Video,Audio,Text Template Instructionalization, GPT Generation
@@
 A-okvqa (Schwenk et al., 2022) Unimodal Understanding Image,Text QA Acc
-POPE (Li et al., 2023b) Unimodal Understanding Image,Text Hallucination Acc
+POPE (Li et al., 2023) Unimodal Understanding Image,Text Hallucination Acc
 IIIT5K (Mishra et al., 2012) Unimodal Understanding Image,Text OCR WAC(word ACC)
@@
 MSRVTT-QA (Xu et al., 2016) Unimodal Understanding Video,Text QA Acc
-MSVD-QA (Xu et al., 2017a) Unimodal Understanding Video,Text QA Acc
+MSVD-QA (Xu et al., 2017) Unimodal Understanding Video,Text QA Acc
 How2QA (Li et al., 2020) Unimodal Understanding Video,Text QA Acc
@@
 STAR (Wu et al., 2021) Unimodal Understanding Video,Text QA Acc
-MSVD-Caption (Xu et al., 2017a) Unimodal Understanding Video,Text QA CIDER
+MSVD-Caption (Xu et al., 2017) Unimodal Understanding Video,Text QA CIDER
 VATEX  (Wang et al., 2019) Unimodal Understanding Video,Text Caption CIDER
@@
 TUT2017 (Mesaros et al., 2016) Unimodal Understanding Audio,Text Classification Acc
-EHSL (Li et al., 2024g) Unimodal Understanding Audio,Text QA Acc
+EHSL (Li et al., 2024f) Unimodal Understanding Audio,Text QA Acc
 Cap3D Caption (Luo et al., 2023b) Unimodal Understanding 3D,Text Caption CIDER
@@
 AVSD (AlAmri et al., 2018) Crossmodal Understanding Video,Audio,Text Dialogue CIDER,BLEU
-RACE-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
+RACE-Audio (Li et al., 2024f) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
 VALOR Caption (Chen et al., 2023e) Crossmodal Understanding Video,Audio,Text Caption CIDER,BLEU
-MMBench-Audio (Li et al., 2024g) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
+MMBench-Audio (Li et al., 2024f) Crossmodal Understanding Image,Audio,Text Comprehensive Benchmark Acc
 AVQA (Li et al., 2022a) Crossmodal Understanding Video,Audio,Text QA Acc
@@
 We statistic the performance of various mainstream Omni-MLLMs in uni-modal understanding, cross-modal understanding, and cross-modal, as shown in Table 5.
-We also show the performance of several Specific-MLLMs (Lin et al., 2024b; Li et al., 2024a; Chu et al., 2023; Xu et al., 2024b; Sun et al., 2024c; Jin et al., 2024a) on selected tasks for comparison.
+We also show the performance of several Specific-MLLMs (Lin et al., 2024; Li et al., 2024a; Chu et al., 2023; Xu et al., 2024b; Sun et al., 2024c; Jin et al., 2024) on selected tasks for comparison.
 The results are mainly from corresponding papers (some results are used as baselines in other papers).
@@
 Table 5: The performance of Omni-MLLMs on different benchmarks.
-The selected uni-modal understanding benchmarks include Video-Text2Text (Xu et al., 2017a, 2016), Image-Text2Text (Goyal et al., 2017; Plummer et al., 2015; Liu et al., 2024d), Audio-Text2Text (Kim et al., 2019; Lipping et al., 2022), and 3D-Text2Text (Deitke et al., 2023).
+The selected uni-modal understanding benchmarks include Video-Text2Text (Xu et al., 2017, 2016), Image-Text2Text (Goyal et al., 2017; Plummer et al., 2015; Liu et al., 2024d), Audio-Text2Text (Kim et al., 2019; Lipping et al., 2022), and 3D-Text2Text (Deitke et al., 2023).
 The chosen uni-modal generation benchmarks include Text2Image (Lin et al., 2014), Text2Video (Xu et al., 2016), and Text2Audio (Kim et al., 2019).